### v1

# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

Abstract
Infrared small target detection poses unique challenges due to the scarcity of intrinsic target features and the abundance of similar background distractors. We argue that background semantics play a pivotal role in distinguishing visually similar objects for this task. To address this, we introduce a new task—-clustered infrared small target detection, and present DenseSIRST, a novel benchmark dataset that provides per-pixel semantic annotations for background regions, enabling the transition from sparse to dense target detection. Leveraging this dataset, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Furthermore, we propose the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which selectively pastes small targets into sky regions during training, avoiding the creation of false alarm targets in complex non-sky backgrounds. Extensive experiments validate the effectiveness of BAG-CP and BAFE-Net in improving target detection accuracy while reducing false alarms. The DenseSIRST dataset, code, and trained models are available at https://github.com/GrokCV/BAFE-Net.
Index Terms:
Infrared Small Target Detection; Clustered Target Detection; Dataset; Feature Exchange; Semantic Segmentation
I Introduction
Figure 1: The small target in the orange box is the real object we are interested in in the sky background, and the small target in the blue box is a virtual alarm in the background of the building. Figure (a) and Figure (c) show clustered small targets against the background of the sky and false alarms against the background of buildings, and figure (b) shows false alarms against the background of buildings.
Infrared imaging has emerged as a crucial tool in various domains, particularly in scenarios where long-distance imaging is required in the absence of visible light, such as night-time operations or environments with limited visibility [1]. By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring [2], maritime target search [3], and rescue missions [4]. However, in these applications, the targets of interest are often located at great distances, resulting in their appearance as extremely small-sized objects in the captured images. Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images, enabling the development of robust and efficient systems that can operate effectively in challenging conditions.
I-A Prior Works on Clustered Infrared Small Target Detection
While the potential of infrared imaging has been recognized across various domains, existing infrared small target detection techniques are predominantly tailored for the detection of sparse targets [5, 6], which proves to be inadequate when faced with complex application scenarios such as drone swarm defense, multi-target tracking systems, and extensive maritime search and rescue operations. Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking dense infrared small targets.
As the distribution of small targets transitions from sparse to dense scenarios, existing infrared small target detection techniques face a multitude of challenges that drastically hinder their effectiveness. Firstly, methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies. Moreover, the reliance on local contrast features [7, 8], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects. Furthermore, even state-of-the-art deep learning methods [9, 10, 11, 12] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, significantly adding to the complexity of detection tasks. Therefore, current methods are plagued by inefficiencies when operating in dense settings, frequently resulting in missed detections and high false alarm rates. This calls for new solutions tailored to overcome the intricacies of dense infrared small target detection.
From another perspective, despite the rapid strides made in generic object detection, with state-of-the-art methods like TOOD [13] and CrossKD [14] addressing dense object detection, these techniques cannot be directly transplanted to the domain of infrared small target detection. The challenges we face in this specialized field stem from several critical factors:
  1. 1.
    Scarcity of Intrinsic Target Features: The foremost challenge stems from the paucity of intrinsic target features. Unlike generic objects, which boast rich visual cues such as appearance, texture, and structure, infrared small targets are minuscule in size, merely popping out as nondescript blobs in the image. This “visual poverty” renders recognition and localization methods based on target-specific characteristics largely ineffectual.
  2. 2.
    Abundance of Similar Background Distractors: Infrared imagery is often plagued by a multitude of background distractors that bear a striking resemblance to the targets of interest. These distractors, such as noise spots and hot spots, also manifest as blobs, making it nearly impossible to discriminate between them and genuine targets based on local visual comparisons.
These challenges inevitably lead us to ponder a fundamental question: What truly distinguishes visually similar objects in single-frame infrared small target detection, assigning them different class labels? Unraveling this enigma holds the key to developing effective and robust algorithms for infrared small target detection, enabling them to navigate the unique challenges posed by this task.
I-B Motivation
In this paper, we posit that the answer to the aforementioned question lies in the background semantics. As illustrated in Fig. 1, visually similar small objects at a distance can be classified differently based on their surrounding context. A small object in the sky background is likely to be a target of interest, while the same object against a building backdrop might be a false alarm. This stark contrast underscores the critical role that background semantics play in assigning meaningful labels to visually similar objects.
However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research. Despite the availability of several public datasets, such as SIRST v1 [15] /v2 [4] , NUDT-SIRST [16], and IRSTD1K [12] , none of these datasets provide detailed semantic annotations for the background regions. To enable the explicit semantic modeling of backgrounds, a prerequisite is the creation of a dataset that includes background class annotations. Such a dataset would provide a solid foundation for researchers to develop algorithms that can leverage the rich contextual information present in the background regions.
To address the aforementioned challenges, this study presents DenseSIRST, a new open-source infrared target detection benchmark dataset. Compared to previous datasets in this field, DenseSIRST has implemented two significant upgrades. Firstly, the dataset transitions from a simple sparse target detection task to the precise detection of dense, clustered targets, substantially increasing the task’s difficulty and practical value. Secondly, DenseSIRST performs detailed per-pixel semantic segmentation of the background, dividing it into sky and non-sky categories, thereby providing preliminary support for explicit semantic modeling of the background.
Building upon this foundation, we propose the Background-Aware Feature Exchange Network (BAFE-Net) framework, which transforms the infrared small target detection paradigm from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a cross-task feature hard-exchange module in the target detection head and background segmentation head, which embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter. To the best of our knowledge, this paper is the first to propose a cross-task feature hard-exchange mechanism. Benefiting from the explicit modeling of background semantics, BAFE-Net can better understand the contextual information in the scene, thereby significantly improving the accuracy of target detection and greatly reducing the false alarm rate.
Finally, by leveraging the fine-grained semantic annotations of the background, we propose a Background-Aware Gaussian Copy-Paste (BAG-CP) method. During the training phase, this method selectively pastes small targets into the sky region based on different background environments during data augmentation. This approach avoids pasting small targets into complex non-sky background regions, such as urban ground, which can produce highly similar false alarm targets.
In summary, our contributions can be categorized into THREE main aspects:
  1. 1.
    We propose a new task—-clustered infrared small target detection, and have released a new dataset, DenseSIRST, annotated with background semantic segmentation.
  2. 2.
    Utilizing DenseSIRST, we propose and demonstrate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which outperforms the original Copy-Paste method that ignores background semantics.
  3. 3.
    We propose the BAFE-Net, which achieves a learning mechanism of mutual promotion between segmentation and detection tasks by implementing feature hard-exchange with a cross-task adapter, based on explicit background modeling.
To validate the effectiveness of BAG-CP and BAFE-Net, we provide extensive experimental results demonstrating their ability to improve the accuracy of target detection while reducing the false alarm rate. The experimental results demonstrate that our method outperforms the compared methods, showcasing its superior performance.
II Related Work
II-A Infrared Small Target Detection
Infrared small target detection methods can be primarily categorized into model-driven and data-driven methods. Model-driven methods can be broadly categorized into three subclasses, each incorporating different strategies to model the relationship between targets and backgrounds.
  1. 1.
    Background Estimation Methods: These include techniques like Top-hat filtering [17] and Max-Median filtering [18]. By leveraging the local consistency of backgrounds, these methods construct filters to enhance target visibility. However, they falter when dealing with extremely weak targets, leading to missed detections and a high sensitivity to hyperparameters. The primary limitation stems from insufficient semantic understanding of the features, which often results in poor performance in complex backgrounds and increased false alarms.
  2. 2.
    Human Visual System (HVS) Methods: Techniques such as local contrast measure (LCM) [7] and multi-scale patch-based contrast measure (MPCM) [19] attempt to emulate the high spatial local contrast sensitivity of the human eye. Despite their intuitive appeal, these methods struggle with very small or blurred targets, mainly due to their inability to generate sufficient contrast.
  3. 3.
    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image model (IPI) [5], partial sum of tensor nuclear norm (PSTNN) [20], and sparse regularization-based twisted tensor model. These methods have gained popularity for their robustness in target and background separation. However, they severely suffer from the challenge of distinguishing background distractors from targets due to the similar sparse characteristic.
Despite their contributions, these methods suffer from several limitations, including poor performance in complex backgrounds, high false alarm rates, and sensitivity to hyperparameters, which we attribute to the lack of feature semantics.
Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging neural networks to automatically extract and learn features relevant for detecting small targets. Methods such as asymmetric context modulation (ACM) [15], dense nested attention networks (DNANet) [16], and gated shaped trans u-net (GSTUnet) [11] have shown significant promise. These methods concentrate on integrating the features of infrared small targets with neural networks, often utilizing deeper neural networks or more robust attention modules. The advantages of data-driven methods lie in their ability to automatically extract features and focus on multi-scale feature interaction, which can be seen as a form of implicit modeling of the relationship between targets and background.
While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances. Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance. Their performance is greatly affected in dense scenes with multiple adjacent targets, where the assumption of local contrast between targets and their neighboring background no longer holds. This fundamental shift in target distribution necessitates the development of approaches that can effectively handle the complex interactions and dependencies among multiple targets in close proximity.
In contrast to previous studies, our work addresses several key limitations and explores new frontiers in infrared small target detection:
  1. 1.
    Focusing on the Unexplored Challenge of Clustered Targets: Unlike traditional approaches that primarily focus on isolated targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field. We further facilitate the research in this area by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
  2. 2.
    Incorporation of Explicit Semantic Modeling with DenseSIRST: In contrast to previous studies that often treat all backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic segmentation annotations for background modeling. Our proposed BAFE-Net thereby gains an enhanced understanding of the target’s background, leading to improved detection accuracy.
  3. 3.
    Introduction of Cross-Task Feature Exchange: Our cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not commonly found in existing work. This approach enables a more precise differentiation between targets and their backgrounds.
II-B Data Augmentation for Object Detection
Data augmentation is pivotal in enhancing the performance and generalization of object detection models. By diversifying training data through various transformations such as flipping, rotation, scaling, and cropping, models learn to recognize and generalize across a broader spectrum of object variations. However, while these basic techniques are widely used, they often fall short in complex detection scenarios, necessitating more sophisticated methods.
To address the limitations of traditional methods, advanced augmentation techniques have been developed, which introduce more nuanced transformations that better preserve and highlight essential features of the objects while still enhancing the dataset’s diversity. For instance, CutMix [21] improves upon the CutOut [22] method by overlaying a segment of one image onto another, blending labels proportionally to the area of the patch, thereby maintaining more contextual integrity. Attentive-CutMix [23] refines this approach by selecting regions based on feature importance, thus preserving significant target details. SnapMix [24] and SaliencyMix [25] adjust the blending process based on saliency or importance maps, aiming to preserve crucial features while enhancing background complexity. TransMix [26] integrates attention mechanisms to better manage the blending of features, focusing on maintaining important target characteristics during augmentation. StyleMix [27] and PuzzleMix [28] introduce style and structural considerations into the mix, allowing for more complex transformations that consider both content and style elements of the images, potentially offering more robust augmentation strategies.
Despite these advancements, these methods primarily assume the presence of larger, easily distinguishable targets and do not specifically cater to the unique challenges posed by clustered infrared small target detection. None of these methods are explicitly designed to handle the high-density and occlusion-prone nature of infrared small targets. This gap highlights the need for augmentation strategies that directly address the intricacies of detecting small, closely-spaced targets against complex backgrounds.
Where our methodology aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and innovative strategies:
  1. 1.
    Pioneering Data Augmentation for Clustered Infrared Small Target Detection: Uniquely, our method pioneers a data augmentation technique specifically tailored for clustered infrared small target detection. It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not often considered in existing techniques.
  2. 2.
    Context-Aware Target Placement: Our proposed BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics. Instead, it integrates background semantics to guide target placement, thereby skirting potential category conflicts induced by inconsistent background contextual semantics. This nuanced approach to the distinctive challenges in infrared small target detection tasks sets our work apart.
II-C Multi-Task Learning and Cross-Task Mechanism
Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning related tasks, such as detection and segmentation. By learning shared representations from multiple supervisory tasks, MTL has shown remarkable success in the field of recognition.
Recent studies have introduced novel architectures and mechanisms to bolster the capabilities of MTL in object detection. Takikawa et al. [29] introduced a two-stream CNN architecture for semantic segmentation that explicitly wires shape information as a separate processing branch, using gates to connect the intermediate layers of the two streams. Misra et al. [30] proposed a new sharing unit called “cross-stitch” unit, which combines activations from multiple networks and can be trained end-to-end to learn an optimal combination of shared and task-specific representations. Zhang et al. [12] developed ISNet, which employs a Taylor finite difference (TFD)-inspired edge block and two-orientation attention aggregation (TOAA) block to enhance edge information and capture shape characteristics of infrared small targets.
However, these methods may not be well-suited for clustered infrared small target detection, as they often rely on edge detection or surface normal estimation, which assume larger targets. In real-world scenarios, infrared small targets typically consist of only 1-9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
In this context, cross-task mechanisms that enable feature exchange between different tasks are more relevant to our work. These methods enhance MTL by facilitating feature exchange or distillation across different tasks, allowing each task to benefit from complementary signals of others. Methods like prediction-and-distillation network (PAD-Net) utilize attention-based mechanisms to distill multi-modal features effectively [31], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [32].
While our work builds upon the foundational principles of multi-task learning and cross-task mechanisms, it distinguishes itself in the feature hard-exchange mechanism. Diverging from the soft fusion approach used in attention-based methods, we employ a feature hard-exchange mechanism for feature fusion. This simple yet effective strategy ensures a more robust integration of features from different tasks.
III The DenseSIRST Dataset
Figure 2: The four images above show diverse backgrounds, including city, sea, cloud, and mountain. The five diagrams below illustrate various targets, including Point, Cluster, UAV, Plane, and Ship.
Figure 3: A total of 12 dataset images are shown. In each set of images, the top figure shows the simulated dense small target, and the bottom figure shows the corresponding sky segmentation of the image.
III-A Motivation and Rationale
The detection of small, low-contrast targets against complex backgrounds in infrared imagery is a significant challenge that impacts the performance of computer vision algorithms. While existing datasets have contributed to the advancement of this field, they often suffer from limitations such as sparse target distribution and a focus on larger target sizes, which do not adequately represent real-world scenarios.
To address these limitations, we introduce the DenseSIRST dataset, a richly annotated collection of infrared images featuring densely populated small targets. DenseSIRST not only increases the number of small targets per image but also encompasses a variety of target sizes and complex clustering, crucial for training and evaluating algorithms under realistic and challenging conditions.
TABLE I: Main characteristics of several popular SIRST datasets.
Datasets Image Type Annotation Type Image Number Target Number Average Target Area Sparse or Clustered Background Semantic Annotation
SIRST [15] Real Pixel 427 533 23 Sparse
SIRSTv2 [4] Real Pixel + BBox + Point 1024 648 24 Sparse
IRSTD1K [12] Real Pixel 1001 1495 38 Sparse
SIRSTAUG [33] Synthetic Pixel 8525 9278 88 Sparse
DenseSIRST (Ours) Real + Synthetic Pixel + BBox + Point 1024 13655 6 Cluster
Table I compares DenseSIRST with several well-regarded datasets, such as SIRST [15], SIRSTv2 [4], IRSTD1K [12], and SIRSTAUG [33]. These datasets predominantly feature sparsely distributed targets and lack comprehensive annotations for clustered small targets. In contrast, DenseSIRST distinguishes itself through:
  1. 1.
    Hybrid Image Composition: Combining real and synthetic targets, DenseSIRST enhances both the realism and diversity of the training data.
  2. 2.
    Target Density and Diversity: With an average target area significantly smaller than existing datasets and a higher number of targets per image, DenseSIRST provides a challenging scenario for detection algorithms.
  3. 3.
    Varied Background Scenarios: The dataset includes images from diverse backgrounds such as urban, mountainous, maritime, and cloudy scenes, as shown in Fig. 2, each contributing to the complexity of the detection tasks.
Fig. 3 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations. These annotations provide valuable contextual information for developing and evaluating detection algorithms.
III-B Target Synthesis and Composition
The DenseSIRST dataset, a blend of real and synthetic infrared images, provides a challenging and dense environment for training robust detection algorithms. We curated this dataset from diverse real-world scenarios, supplemented with synthetic targets designed for physical plausibility and visual realism.
The dataset leverages the existing SIRSTv2, creating a broad library of infrared small targets. Each image is meticulously designed to emulate the densely populated surveillance scenarios. Specifically, one to three “dense areas” are defined randomly within the less cluttered sky regions of each base image. These areas, each measuring $20\times 20$ pixels, are populated with 8 to 12 targets selected randomly from the target library. The selected targets are resized to dimensions within $5\times 5$ pixels to maintain scale variability and pasted with a spacing of 1 to 2 pixels between them to simulate realistic clustering.
To ensure a smooth integration of targets and their backgrounds, we employ a Gaussian-weighted copy-paste technique, an evolution of the concept from mixup. This modification is governed by the following equation:
$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1]$ (1)
Here, $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location. Different from mixup, its weight is set to $1$ for enhancing the reality of the synthesized target. $B_{box}$ represents the original target appearance. The scalar $\lambda$, which adjusts the target’s brightness, is randomized to diversify the brightness and contrast of targets, thus enriching the sample space.
The matrix $G$ denotes a Gaussian blur, employed during target pasting, and is defined as:
$G(x,y)=e^{-\frac{{(x-\rho_{x}w)^{2}}}{{2\sigma_{x}^{2}}}-\frac{{(y-\rho_{y}h)^{2}}}{{2\sigma_{y}^{2}}}},$ (2)
The relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$. These values, when multiplied by the target bounding box’s width $w$ and height $h$, yield the actual pixel offsets, ensuring appropriate scaling with the target size. The Gaussian distribution’s spread is controlled by the standard deviations $\sigma_{x}$ and $\sigma_{y}$ along the $x$ and $y$ axes, randomized within $[0.3,0.6]$. This randomness guarantees varying degrees of edge blurring among the simulated targets, reinforcing their realistic appearance.
To further enhance variability, each target undergoes a rotational transformation. The Gaussian matrix’s rotation $\theta$ is applied as:
$\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x-\mu_{x}\\ y-\mu_{y}\end{bmatrix},$ (3)
Here, $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
III-C Dataset Description and Characteristics
III-C1 Composition of the Dataset
The DenseSIRST dataset comprises 1024 infrared images, with a total of 13,655 densely clustered small targets. The images in the DenseSIRST dataset cover a wide range of realistic scenarios, including urban areas, mountainous regions, maritime environments, and cloudy scenes. This diversity enables the development and assessment of infrared small target detection algorithms that can generalize well to various real-world applications. The division of training set, validation set, and test set is aligned with the SIRSTv2 dataset.
III-C2 Annotations and Labels
In the DenseSIRST dataset, each image is meticulously annotated with three types of labels: pixel-level masks, bounding boxes, and point annotations. The pixel-level masks provide precise segmentation of each target, enabling the evaluation of algorithms at a fine-grained level. Bounding box annotations encapsulate each target with a rectangular box, facilitating the assessment of detection performance using metrics such as precision, recall, and F1 score. Point annotations mark the center of each target, allowing for the evaluation of localization accuracy.
III-C3 Statistical Characteristics of the Dataset
The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection. This part offers a detailed analysis of the dataset’s key characteristics, including the distribution of image sizes, target sizes, local contrast, and brightness levels.
Figure 4: Figure (a) is the image resolution characteristic, the larger the range of circles in the scatter plot, the greater the number of images at that resolution. Figure (b) is the target size characteristic, and the larger the range of circles in the scatter plot, the greater the number of small targets of that size. Figure (c) shows the percentage of different local contrasts for small objects in the dataset, with the lowest local contrast being 0.15. Figure (d) shows the percentage of brightness distribution for small targets in the dataset, and shows that the brightness distribution is wide from dark targets to light targets.
Image Size Distribution: The dataset, as depicted in Fig. 4 (a), includes images with resolutions ranging from $320\times 256$ to $1280\times 1024$ pixels, covering various typical infrared camera resolutions encountered in real-world scenarios. This diversity is crucial for developing detection algorithms that can generalize across different resolutions.
Target Size Distribution: Most targets, as indicated in Fig. 4 (b), are extremely small, predominantly below $5\times 5$ pixels. The $3\times 3$ pixels size is the most frequent, aligning with the characteristics of small target detection in infrared imagery. With a total of 13,655 annotated targets, the dataset provides a rich set of examples for training and evaluation.
Local Contrast Distribution: The vast majority (90%) of targets have local contrast values below $2$, as shown in Fig. 4 (c). This low contrast is indicative of the inherent difficulty in detecting such targets against cluttered backgrounds. Moreover, the presence of nearby targets within dense clusters can inadvertently be included in the background calculation, leading to an underestimation of local contrast and further complicating detection.
Target Brightness Distribution: Fig. 4 (d) reveals that only a minority of targets are the brightest points in their respective images. This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection in infrared images.
In conclusion, the DenseSIRST dataset, with its wide range of image dimensions, prevalence of small target sizes, and low local contrast, provides a realistic and challenging benchmark for infrared small target detection algorithms.
IV Background-Aware Feature Exchange Network
IV-A Overall Architecture and Training Pipeline
Figure 5: Schematic diagram of the proposed BAFE-Net network. BAFE-Net contains three modules: ResNet, FPN and BAFE-Head.
Figure 6: First, a dense area is randomly selected from the sky area of the image, and then several small targets are randomly selected from the small target library. After determining the appropriate pasting position, the Gaussian technique is used to integrate the small targets into the new background.
The proposed BAFE-Net adopts a fully convolutional one-stage object detector (FCOS) [34] as its host network. FCOS is a simple yet effective anchor-free object detection framework that has demonstrated superior performance in various computer vision tasks. By leveraging the inherent advantages of FCOS, such as its ability to handle objects of different scales and aspect ratios without the need for predefined anchor boxes, BAFE-Net aims to address the challenges of detecting small, densely clustered targets in infrared imagery.
The overall architecture of BAFE-Net is illustrated in Fig. 5. In addition to the original detection head present in the baseline FCOS, BAFE-Net introduces a segmentation head within the BAFE-Head module. The primary purpose of incorporating this segmentation head is to enable explicit modeling of background semantics, which is crucial for enhancing the network’s ability to distinguish between real targets and background clutter in infrared scenes. By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information, ultimately leading to improved detection performance.
In addition to the architecture modifications, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP). This technique is used during the training phase to augment the training dataset with realistically integrated synthetic targets. BAG-CP strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, which significantly enhances the training process. The use of Gaussian smoothing at the edges of the pasted targets ensures seamless integration, avoiding artificial boundary artifacts that could mislead the learning algorithm.
This context-aware synthesis and integration approach allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds. By training the network to recognize and handle these challenging conditions effectively, BAFE-Net sets a new standard for performance in infrared small target detection.
IV-B Cross-Task Feature Exchange Module
During the feature extraction process, partial feature exchange is performed between segmentation and classification tasks within this network. The segmentation and classification tasks have different focuses when learning features: segmentation emphasizes local details and contextual information, whereas classification emphasizes global features. Despite their differing objectives, there is a potential synergistic relationship between segmentation and classification tasks. By exchanging features, information can be shared between the two tasks, promoting collaborative learning and enhancing the feature representation capabilities of each task. Segmentation can assist classification in better understanding the structure and semantics of the image, while classification can help segmentation in distinguishing boundaries and details of different categories. Specifically, segmentation features can leverage global information from classification features to improve overall understanding, while classification features can utilize local detail information from segmentation features to enhance precision. Mutual learning between segmentation and classification features can reduce the risk of overfitting and improve the model’s performance in practical applications. Feature exchange enables the model to perform more robustly when facing different tasks. It allows the model to better adapt to various changes and uncertainties, thereby enhancing its generalization ability across different scenarios.
Feature exchange has been proven feasible for several reasons: 1) Complementarity of features: Although segmentation and classification features are different, they are extracted from the same input image and thus possess a degree of complementarity. By exchanging features, each task can utilize useful information from the other task, thereby improving overall performance. 2) Compatibility of features: Deep learning models use continuous and differentiable operations, such as convolution and pooling, when extracting features. Through feature exchange, these operations can still ensure the coherence and differentiability of features, thus not disrupting the training process of the model. 3) Support from existing research: Some studies have already shown that in multi-task learning, sharing or exchanging features can improve the overall performance of tasks. For example, in the Changer [35] paper, channel and spatial exchanges were used and proven feasible.
Feature exchange primarily consists of two methods: channel exchange and spatial exchange. Channel exchange involves directly swapping features at the channel level between different feature maps, while spatial exchange is achieved by swapping features along the spatial dimensions of the feature maps. In the Cross-Task Feature Exchange Module, channel exchange is employed, and experiments have shown that channel exchange performs better than spatial exchange. The experimental results will be presented in Section V.
Assume ${x}_{1}$ and ${x}_{2}$ are the input tensors, each with a shape of $(N,C,H,W)$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width, respectively. Let $p$ be a parameter controlling the proportion of channel exchange. Define an exchange mask $E$ with a shape of $C$,
${E}[i]=\begin{cases}1&\text{, if }i\mod\left(\frac{1}{p}\right)=0\\ 0\end{cases}$ (4)
where $i$ is the channel index. Expand the exchange mask $E$ to the shape of $(N,C,H,W)$,
${E}_{\text{expanded}}={E}\otimes{1}_{N\times H\times W}$ (5)
where $\otimes$ denotes the expansion operation, and ${1}_{N\times H\times W}$ is a tensor of ones with a shape of $(N,H,W)$. Define the output tensors ${y}_{1}$ and ${y}_{2}$ ,
${y}_{1}={x}_{1}\odot(1-{E}_{\text{expanded}})+{x}_{2}\odot{E}_{\text{expanded}}$ (6)
${y}_{2}={x}_{2}\odot(1-{E}_{\text{expanded}})+{x}_{1}\odot{E}_{\text{expanded}}$ (7)
where $\odot$ denotes element-wise multiplication.
However, the features to be exchanged are cross-modal. Therefore, an Adapter is introduced to process these features before the exchange. Here, an MLP (Multi-Layer Perceptron) is used as the Adapter. The features to be exchanged are fed into the Adapter separately, and after processing, the features are exchanged. This approach addresses the cross-modal issue between the features.
IV-C Background-Aware Gaussian Copy-Paste
Infrared small target detection is inherently challenged by the paucity of intrinsic visual features in targets and the interference from similar background elements. Traditional data augmentation strategies, such as random cropping or flipping, are not effective in addressing these unique challenges. Consequently, we propose a novel data augmentation technique, Background-Aware Gaussian Copy-Paste (BAG-CP), which capitalizes on contextual information to enhance the training process. BAG-CP selectively augments training images by pasting infrared small targets into semantically relevant background areas, generating more realistic and diverse training samples. The fundamental idea is to bolster the detector’s ability to discern real targets from background noise through the seamless blending of pasted targets into their new environment. As illustrated in Fig. 6, The BAG-CP process unfolds as follows:
Prerequisite: Semantic Annotation of Background: Before performing the copy-paste operation, it is crucial to have the image segmented into different semantic regions. The DenseSIRST dataset includes annotated background categories, providing detailed background masks that separate sky from non-sky regions.
Step 1: Selection and Extraction of Infrared Small Targets: A library of small targets is created during the generation of the DenseSIRST dataset. A random selection of small targets from this library is made for pasting into the current training image.
Step 2: Gaussian Smoothing of Edges: Prior to pasting, the selected targets are subjected to Gaussian smoothing. This step ensures a smooth transition around the pasted targets, minimizing any conspicuous artifacts that might disrupt the learning process.
Step 3: Context-Aware Target Pasting Based on Background Semantics: The core of BAG-CP lies in context-aware placement. Using the semantic annotation of the background, we classify it into sky and non-sky regions. When choosing where to paste the Gaussian-smoothed targets, we only select locations within the sky regions. This approach mimics real-world scenarios where infrared small targets, especially aerial targets, are more likely to appear against a sky background.
During the pasting process, 1-3 regions are selected within the image to form clusters of small targets. Each region is 20 × 20 pixels in size and is populated with 8 to 12 targets randomly chosen from the target library. The selected targets are resized to within 5 × 5 pixels to maintain proportional variability and are pasted with a spacing of 1 to 2 pixels to simulate realistic clustering. The Gaussian-weighted edge blurring technique used follows Eq. (1), Eq. (2), and Eq. (3).
V Experiments
In this section, we outline a series of experimental evaluations designed to validate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method and the Background-Aware Feature Exchange Network (BAFE-Net) proposed in this study. Additionally, we conduct comparative analyses with state-of-the-art methods in infrared small target detection, using both quantitative metrics and qualitative assessments. We address the following pivotal questions in our experimental analysis:
  1. 1.
    Ablation Study on Background-Aware Gaussian Copy-Paste. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
  2. 2.
    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlights the contributions of each component in BAFE-Net.
  3. 3.
    Feature Exchange Mechanism Analysis. We investigate the impact of different feature exchange strategies on the performance of BAFE-Net to determine the optimal approach for exchanging features between the target detection head and the background segmentation head.
  4. 4.
    Impact of Backbone Depth: We examine the influence of varying backbone depths on the performance of the BAFE-Net. The primary aim is to establish the optimal depth that maximizes target detection while balancing computational efficiency.
  5. 5.
    Comparative Analysis with State-of-the-Art Methods. We conduct a comprehensive comparison between our method and other state-of-the-art methods on the DenseSIRST dataset.
V-A Experimental Settings
1) Dataset: During the experimental phase of our research, we conducted tests on the proposed dataset DenseSIRST.
2) Implementation Details: Our network was trained on an NVIDIA RTX 3090 GPU platform with 24GB of memory. In our experiments, we employed the DeepInfrared toolkit for object detection. Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion. We choose FocalLoss as the loss function of the classification branch, IoULoss as the loss function of the regression branch, and CrossEntropyLoss as the loss function of the segmentation branch. We choose DAdaptAdam as the optimizer, lr is set to 1.0, and weight decay is set to 0.05. Considering training and convergence time, we trained for 20 epochs within the DeepInfrared framework. Through experimentation, we observed that different parameter settings affect the detection efficacy of dense small objects. Thus, while ensuring basic fairness, we adjusted the parameters of different methods to obtain optimal detection results.
3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment. They are defined as:
$\text{mAP}=\frac{1}{N}\sum_{i=1}^{N}\text{AP}(i),$ (8)
[…]
$\text{Recall}=\frac{TP}{TP+FN},$ (10)
Here, $N$ is the total number of samples, $TP$, $FP$, and $FN$ denote the number of true positive, false positive, and false negative instances, respectively. Precision $p(r)$ is defined as the proportion of true positive detections out of the total number of positive detections at a given recall level $r$.
TABLE II: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net
Strategy Module DenseSIRST
Copy-Paste Cluster Gaussian Background-Aware mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 0.225 0.311 0.180 0.314
(b) $\checkmark$ 0.236 0.322 0.203 0.303
(c) $\checkmark$ $\checkmark$ 0.242 0.329 0.205 0.335
(d) $\checkmark$ $\checkmark$ 0.245 0.321 0.211 0.338
(e) $\checkmark$ $\checkmark$ $\checkmark$ 0.250 0.319 0.215 0.315
(f) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.258 0.340 0.220 0.340
TABLE III: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
Strategy Module DenseSIRST
Seg Head Feature Exchange Adapter mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 0.225 0.311 0.180 0.314
(b) $\checkmark$ 0.246 0.324 0.183 0.304
(c) $\checkmark$ $\checkmark$ 0.254 0.321 0.199 0.310
(d) $\checkmark$ $\checkmark$ $\checkmark$ 0.262 0.318 0.208 0.305
TABLE IV: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
Strategy Hidden layers DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 8 0.222 0.305 0.206 0.308
(b) 16 0.243 0.318 0.214 0.310
(c) 32 0.256 0.329 0.218 0.312
(d) 64 0.266 0.336 0.230 0.323
TABLE V: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
Strategy Exchange Method DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) Channel 0.266 0.336 0.230 0.323
(b) Spatial 0.223 0.304 0.218 0.322
(c) Channel + Spatial 0.246 0.310 0.199 0.314
(d) Spatial + Channel 0.224 0.310 0.201 0.328
TABLE VI: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
Strategy ResNet DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) ResNet18 0.258 0.328 0.233 0.341
(b) ResNet34 0.258 0.330 0.226 0.332
(c) ResNet50 0.270 0.332 0.236 0.329
(d) ResNet101 0.256 0.324 0.241 0.336
TABLE VII: Comparison with Other State-of-the-art methods on DenseSIRST.
Method Backbone DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ Flops $\downarrow$ Params $\downarrow$
              One-stage
FCOS [34] ResNet50 0.232 0.315 0.204 0.324 50.291G 32.113M
SSD [36] 0.211 0.421 0.178 0.424 87.552G 23.746M
GFL [37] ResNet50 0.253 0.332 0.230 0.317 52.296G 32.258M
ATSS [38] ResNet50 0.248 0.327 0.202 0.326 51.504G 32.113M
CenterNet [39] ResNet50 0.000 0.000 0.000 0.000 50.278G 32.111M
PAA [40] ResNet50 0.255 0.545 0.228 0.551 51.504G 32.113M
PVT-T [41] 0.109 0.481 0.093 0.501 41.623G 21.325M
RetinaNet [42] ResNet50 0.114 0.510 0.086 0.523 52.203G 36.330M
EfficientDet [43] 0.099 0.433 0.072 0.419 34.686G 18.320M
TOOD [13] ResNet50 0.256 0.355 0.226 0.342 50.456G 32.018M
VFNet [44] ResNet50 0.253 0.336 0.214 0.336 48.317G 32.709M
YOLOF [45] ResNet50 0.091 0.009 0.002 0.009 25.076G 42.339M
AutoAssign [46] ResNet50 0.255 0.354 0.180 0.314 50.555G 36.244M
DyHead [47] ResNet50 0.249 0.335 0.189 0.328 27.866G 38.890M
Two-stage
Faster R-CNN [48] ResNet50 0.091 0.022 0.015 0.029 0.759T 33.035M
Cascade R-CNN [49] ResNet50 0.136 0.188 0.139 0.194 90.978G 69.152M
Dynamic R-CNN [50] ResNet50 0.184 0.235 0.111 0.190 63.179G 41.348M
Grid R-CNN [51] ResNet50 0.091 0.018 0.025 0.037 0.177T 64.467M
Libra R-CNN [52] ResNet50 0.141 0.142 0.085 0.120 63.990G 41.611M
End2End
DETR [53] ResNet50 0.000 0.000 0.000 0.000 24.940G 41.555M
Deformable DETR [54] ResNet50 0.024 0.016 0.018 0.197 51.772G 40.099M
DAB-DETR [55] ResNet50 0.005 0.054 0.000 0.001 28.939G 43.702M
Conditional DETR [56] ResNet50 0.000 0.000 0.000 0.001 27.143G 40.297M
Sparse R-CNN [57] ResNet50 0.183 0.572 0.154 0.614 45.274G 0.106G
$\star$ BAFE-Net (Ours) ResNet50 0.270 0.332 0.236 0.329 69.114G 35.329M
V-B Ablation Study
V-B1 Ablation Study on Background-Aware Gaussian Copy-Paste
In this study, we initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting. Our aim is to empirically validate the enhancements to detection performance brought about by these steps. To achieve this, we explore various combinations of the aforementioned components and conduct experimental validation of the results.
The experimental outcomes, as outlined in Table II, generate several key insights, providing a holistic understanding of the impact of each module. Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study. The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
Further refinement through clustering of pasted objects also yields substantial improvements. This is evident when comparing strategies (b) to (c), and (d) to (e), respectively, indicating the considerable value of object clustering. Applying Gaussian blur to the edges of the pasted objects, as a means to enhance their realistic appearance, also proves to be significantly beneficial in improving detection performance. This is substantiated when comparing strategies (b) with (d), and (c) with (e).
Finally, the culmination of all modules in a Background-Aware Gaussian Copy-Paste context, as depicted by strategy (f), results in further enhancements in detection capability. This is demonstrated with an additional increase of 0.8% on the DenseSIRST dataset when comparing strategy (e) with strategy (f).
V-B2 Ablation Study on the Scheme of Background-Aware Feature Exchange
The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table III. To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
Table III encapsulates the results of our experimental study and yields several crucial insights. Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research. Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head. This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
Further, juxtaposition of Strategy (b) and Strategy (c) highlights the contribution of the Feature Exchange in the interaction mode between the detection head and the segmentation head. This interaction led to a notable improvement in detection performance by 1.1% on the DenseSIRST dataset. Upon comparing Strategy (c) and Strategy (d), it becomes evident that introducing feature adapation prior to the Feature Exchange further amplifies detection performance.
As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration. Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity. To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  IV. The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance. However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially. Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
V-B3 Ablation Study on Feature Exchange Mechanism
In this part, a rigorous exploration of various Feature Exchange Mechanisms was undertaken. This encompassed diverse mechanisms such as Channel Exchange and Spatial Exchange, along with a multitude of their combinations. The quest was to ascertain the most efficacious method for Feature Exchange. Channel Exchange, as a mechanism, facilitates the direct interplay at the channel level amongst different features. In contrast, Spatial Exchange is implemented by orchestrating exchanges along the spatial dimensions inherent in the feature maps. The exchange ratio was judiciously set at 0.5 to maintain a balance.
As delineated in Table V, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested. The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
V-B4 Impact of Backbone Depth:
For infrared small target detection, the depth of the backbone network emerges as a critical factor. A shallower network, while computationally efficient, might extract features with insufficient semantic richness, thereby impinging on the overall detection performance. Conversely, an excessively deep network might overwhelm the delicate features of smaller targets amidst the background, consequently leading to a performance degradation.
We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net. The results of these experiments are summarized in Table VI. Notably, ResNet50 emerged as the most suitable backbone for the BAFE-Net.
V-C Comparison with State-of-the-Arts
We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall. The experimental results are presented in Table VII.
End-to-end (End2End) methods, while theoretically appealing due to their NMS-free style, demonstrated suboptimal performance. This can be attributed to their inherent requirement for extensive training data, akin to COCO-level volumes, for effective operation. The relatively limited availability of infrared small target detection datasets inherently handicaps such methods, resulting in their diminished performance.
Next, we observed that two-stage methods underperformed compared to one-stage methods. Two-stage methods generate candidate regions based on deeper features, which are then subjected to classification and bounding box regression. However, the DenseSIRST dataset, characterized by small and densely clustered infrared targets, poses a challenge for such methods, leading to their suboptimal performance.
In contrast, one-stage methods, which directly regress bounding boxes and categories on the feature pyramids, demonstrated superior performance. These methods are not only faster and more computationally efficient, but also achieved higher mAP and recall, indicating their overall effectiveness. Significantly, our BAFE-Net method outperformed all other one-stage methods. The superior performance of BAFE-Net validates the efficacy of our approach in accurately detecting small infrared targets amidst dense conditions.
V-D Visual Analysis
To further validate the effectiveness of our proposed methods, we conduct a visual analysis by comparing the feature activation maps of our approach with those of the top-performing competing methods in the quantitative evaluation. Fig. 7 illustrates the feature activation maps of the five most competitive methods, including our own, on representative samples from the DenseSIRST dataset.
From the visual results, we observe that our method exhibits two key advantages over the other approaches. First, the feature activation maps of our method demonstrate significantly stronger activation intensities in the target regions, indicating a higher sensitivity to the presence of infrared small targets. The deeper activation levels in the target areas suggest that our method is more effective in capturing the discriminative features of the targets, contributing to improved detection accuracy.
Second, our method generates cleaner background activation maps compared to the other methods. As evident from Fig. 7, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference. This can be attributed to the background-aware mechanism employed by BAFE-Net, which learns to differentiate between target and background regions during the feature exchange process. By minimizing the activation levels in the background areas, our method demonstrates enhanced robustness against false positives, leading to a higher detection precision.
Figure 7: Feature visualization comparison. Compared with other methods, the proposed BAFE-Net shows superior performance with higher detection of small objects and fewer false alarms.
VI Conclusion
This study tackles the key challenges in infrared small target detection by leveraging background semantics. We introduce DenseSIRST, a benchmark dataset with detailed background annotations, enabling the transition from sparse to dense target detection. Building upon this, we propose BAFE-Net, a multi-task architecture that explicitly models background semantics and facilitates cross-task feature hard-exchange. Moreover, we develop BAG-CP, a data augmentation method that selectively pastes targets into sky regions, outperforming the original Copy-Paste method. Extensive experiments validate the effectiveness of BAFE-Net and BAG-CP in improving detection accuracy and reducing false alarms. Our work underscores the importance of explicitly leveraging contextual information in infrared small target detection and paves the way for the development of more robust and context-aware algorithms in this domain.

### v2

@@
 Leveraging this dataset, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation.
-BAFE-Net introduces a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks.
+BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks.
 Furthermore, we propose the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which selectively pastes small targets into sky regions during training, avoiding the creation of false alarm targets in complex non-sky backgrounds.
@@
 I Introduction
-Figure 1: The small target in the orange box is the real object we are interested in in the sky background, and the small target in the blue box is a virtual alarm in the background of the building.
-Figure (a) and Figure (c) show clustered small targets against the background of the sky and false alarms against the background of buildings, and figure (b) shows false alarms against the background of buildings.
+Figure 1: Illustration of the background semantics’ crucial role in infrared small target detection.
+The small target in the orange box represents a genuine target of interest against the sky background, while the small target in the blue box is a false alarm in the context of the building background.
+When considering only local regions, it is challenging to differentiate between clustered small targets and false alarms due to their similar appearances.
+By incorporating global contextual information, the distinction between real targets and false alarms becomes more apparent.
 Infrared imaging has emerged as a crucial tool in various domains, particularly in scenarios where long-distance imaging is required in the absence of visible light, such as night-time operations or environments with limited visibility [1].
-By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring [2], maritime target search [3], and rescue missions [4].
+By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring, maritime target search [2], and rescue missions [3].
 However, in these applications, the targets of interest are often located at great distances, resulting in their appearance as extremely small-sized objects in the captured images.
-Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images, enabling the development of robust and efficient systems that can operate effectively in challenging conditions.
+Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images.
 I-A Prior Works on Clustered Infrared Small Target Detection
 While the potential of infrared imaging has been recognized across various domains, existing infrared small target detection techniques are predominantly tailored for the detection of sparse targets [4, 5], which proves to be inadequate when faced with complex application scenarios such as drone swarm defense, multi-target tracking systems, and extensive maritime search and rescue operations.
-Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking dense infrared small targets.
+Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking clustered infrared small targets.
 As the distribution of small targets transitions from sparse to dense scenarios, existing infrared small target detection techniques face a multitude of challenges that drastically hinder their effectiveness.
 Firstly, methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies.
-Moreover, the reliance on local contrast features [7, 8], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects.
+Moreover, the reliance on local contrast features [6], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects.
 Furthermore, even state-of-the-art deep learning methods [7, 8, 9, 10] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, significantly adding to the complexity of detection tasks.
+For instance, Dai et al.
+[11] introduced a detection transformer designed for arbitrarily oriented objects, yet its applicability to dense infrared small target settings remains limited due to the lack of intrinsic target features.
 Therefore, current methods are plagued by inefficiencies when operating in dense settings, frequently resulting in missed detections and high false alarm rates.
 This calls for new solutions tailored to overcome the intricacies of dense infrared small target detection.
-From another perspective, despite the rapid strides made in generic object detection, with state-of-the-art methods like TOOD [13] and CrossKD [14] addressing dense object detection, these techniques cannot be directly transplanted to the domain of infrared small target detection.
-The challenges we face in this specialized field stem from several critical factors:
+While the rapid advancements in generic object detection have led to state-of-the-art methods like TOOD [12] and CrossKD [13] that address dense object detection, these techniques cannot be directly applied to the unique challenges posed by infrared small target detection.
+The specialized nature of this field presents several critical factors that hinder the direct transplantation of generic vision approaches:
   1. 1.
@@
 These distractors, such as noise spots and hot spots, also manifest as blobs, making it nearly impossible to discriminate between them and genuine targets based on local visual comparisons.
+Liao et al.
+[14] tackled similar challenges in RGB-thermal salient object detection, where distortions like noise and motion blur compounded the complexity of separating salient objects from background clutter.
 These challenges inevitably lead us to ponder a fundamental question: What truly distinguishes visually similar objects in single-frame infrared small target detection, assigning them different class labels?
-Unraveling this enigma holds the key to developing effective and robust algorithms for infrared small target detection, enabling them to navigate the unique challenges posed by this task.
+Unraveling this enigma holds the key to developing effective and robust algorithms for this task.
 I-B Motivation
@@
 However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research.
-Despite the availability of several public datasets, such as SIRST v1 [15] /v2 [4] , NUDT-SIRST [16], and IRSTD1K [12] , none of these datasets provide detailed semantic annotations for the background regions.
+Despite the availability of several public datasets [15, 3, 16, 10], none of these datasets provide detailed semantic annotations for the background regions.
 To enable the explicit semantic modeling of backgrounds, a prerequisite is the creation of a dataset that includes background class annotations.
@@
 Building upon this foundation, we propose the Background-Aware Feature Exchange Network (BAFE-Net) framework, which transforms the infrared small target detection paradigm from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation.
-BAFE-Net introduces a cross-task feature hard-exchange module in the target detection head and background segmentation head, which embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter.
+BAFE-Net introduces a dynamic cross-task feature hard-exchange module in the target detection head and background segmentation head, which selectively embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter.
 To the best of our knowledge, this paper is the first to propose a cross-task feature hard-exchange mechanism.
@@
   3. 3.
-    We propose the BAFE-Net, which achieves a learning mechanism of mutual promotion between segmentation and detection tasks by implementing feature hard-exchange with a cross-task adapter, based on explicit background modeling.
-To validate the effectiveness of BAG-CP and BAFE-Net, we provide extensive experimental results demonstrating their ability to improve the accuracy of target detection while reducing the false alarm rate.
-The experimental results demonstrate that our method outperforms the compared methods, showcasing its superior performance.
+    We propose the BAFE-Net, which achieves a dynamic mechanism of mutual promotion between segmentation and detection tasks by implementing selective feature hard-exchange with a cross-task adapter, based on explicit background modeling.
 II Related Work
 II-A Infrared Small Target Detection
-Infrared small target detection methods can be primarily categorized into model-driven and data-driven methods.
-Model-driven methods can be broadly categorized into three subclasses, each incorporating different strategies to model the relationship between targets and backgrounds.
+Infrared small target detection methods can be broadly classified into two main categories: model-driven and data-driven approaches.
+The model-driven methods, in particular, encompass three distinct subclasses, each employing unique strategies to capture and represent the relationship between targets and their surrounding backgrounds.
   1. 1.
-    Background Estimation Methods: These include techniques like Top-hat filtering [17] and Max-Median filtering [18].
-By leveraging the local consistency of backgrounds, these methods construct filters to enhance target visibility.
+    Background Estimation Methods: By leveraging the local consistency of backgrounds, these methods construct filters to enhance target saliency.
 However, they falter when dealing with extremely weak targets, leading to missed detections and a high sensitivity to hyperparameters.
-The primary limitation stems from insufficient semantic understanding of the features, which often results in poor performance in complex backgrounds and increased false alarms.
+The limitation stems from insufficient semantic understanding of the features [17], which results in poor performance in complex backgrounds and increased false alarms.
   2. 2.
@@
   3. 3.
-    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image model (IPI) [5], partial sum of tensor nuclear norm (PSTNN) [20], and sparse regularization-based twisted tensor model.
-These methods have gained popularity for their robustness in target and background separation.
+    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image (IPI) model [4] and reweighted infrared patch-tensor (RIPT) model [5].
+These methods have gained popularity due to their superior background suppression capabilities compared to previous techniques.
 However, they severely suffer from the challenge of distinguishing background distractors from targets due to the similar sparse characteristic.
 Despite their contributions, these methods suffer from several limitations, including poor performance in complex backgrounds, high false alarm rates, and sensitivity to hyperparameters, which we attribute to the lack of feature semantics.
-Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging neural networks to automatically extract and learn features relevant for detecting small targets.
+Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging deep networks to automatically learn features relevant for detecting small targets.
+Recent advancements have also focused on implementing these deep learning-based methods efficiently, such as the work on high-accuracy and real-time two-stage small object detection on FPGA [19].
 Methods such as asymmetric context modulation (ACM) [15], dense nested attention networks (DNANet) [16], and gated shaped trans U-Net (GSTUnet) [9] have shown significant promise.
-These methods concentrate on integrating the features of infrared small targets with neural networks, often utilizing deeper neural networks or more robust attention modules.
+These methods concentrate on integrating the features of infrared small targets with neural networks, utilizing deeper neural networks or more robust attention modules.
+Similar advancements in applying deep learning to infrared imaging have been seen in related fields, such as the work on infrared-guided eye localization for autostereoscopic displays [20].
 The advantages of data-driven methods lie in their ability to automatically extract features and focus on multi-scale feature interaction, which can be seen as a form of implicit modeling of the relationship between targets and background.
+Similar approaches have been applied in other domains, such as cloud detection, where multi-level information fusion with edge information injection has been used to improve detection performance [21].
+In a related field, transformer-based models have been employed for spatio-temporal unsupervised traffic anomaly detection in aerial videos [22], demonstrating the versatility of these advanced deep learning techniques in handling complex visual data with small objects or anomalies.
 While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances.
-Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance.
+Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance [23].
 Their performance is greatly affected in dense scenes with multiple adjacent targets, where the assumption of local contrast between targets and their neighboring background no longer holds.
 This fundamental shift in target distribution necessitates the development of approaches that can effectively handle the complex interactions and dependencies among multiple targets in close proximity.
-In contrast to previous studies, our work addresses several key limitations and explores new frontiers in infrared small target detection:
+In contrast to previous studies, our work explores a new detection paradigm for infrared small targets:
   1. 1.
-    Focusing on the Unexplored Challenge of Clustered Targets: Unlike traditional approaches that primarily focus on isolated targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field.
-We further facilitate the research in this area by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
+    Unexplored Challenge for Clustered Targets: Unlike traditional approaches that focus on sparse targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field.
+We further facilitate the research by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
   2. 2.
-    Incorporation of Explicit Semantic Modeling with DenseSIRST: In contrast to previous studies that often treat all backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic segmentation annotations for background modeling.
+    Explicit Semantic Modeling: In contrast to previous studies that treat backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic annotations for background modeling.
 Our proposed BAFE-Net thereby gains an enhanced understanding of the target’s background, leading to improved detection accuracy.
   3. 3.
-    Introduction of Cross-Task Feature Exchange: Our cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not commonly found in existing work.
+    Introduction of Cross-Task Feature Exchange: Our dynamic cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not found in existing work.
 This approach enables a more precise differentiation between targets and their backgrounds.
@@
 This gap highlights the need for augmentation strategies that directly address the intricacies of detecting small, closely-spaced targets against complex backgrounds.
-Where our methodology aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and innovative strategies:
+Although our BAG-CP aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and strategies:
   1. 1.
     Pioneering Data Augmentation for Clustered Infrared Small Target Detection: Uniquely, our method pioneers a data augmentation technique specifically tailored for clustered infrared small target detection.
-It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not often considered in existing techniques.
+It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not considered in existing techniques.
   2. 2.
-    Context-Aware Target Placement: Our proposed BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics.
+    Background-Aware Target Placement: Our BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics.
 Instead, it integrates background semantics to guide target placement, thereby skirting potential category conflicts induced by inconsistent background contextual semantics.
-This nuanced approach to the distinctive challenges in infrared small target detection tasks sets our work apart.
 II-C Multi-Task Learning and Cross-Task Mechanism
-Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning related tasks, such as detection and segmentation.
-By learning shared representations from multiple supervisory tasks, MTL has shown remarkable success in the field of recognition.
-Recent studies have introduced novel architectures and mechanisms to bolster the capabilities of MTL in object detection.
+Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning shared representations from related tasks, such as detection and segmentation.
 Takikawa et al.
@@
 Zhang et al.
-[12] developed ISNet, which employs a Taylor finite difference (TFD)-inspired edge block and two-orientation attention aggregation (TOAA) block to enhance edge information and capture shape characteristics of infrared small targets.
+[10] developed ISNet, which employs a Taylor finite difference inspired edge block and two-orientation attention aggregation block to enhance edge information and capture shape characteristics of infrared small targets.
 However, these methods may not be well-suited for clustered infrared small target detection, as they often rely on edge detection or surface normal estimation, which assume larger targets.
-In real-world scenarios, infrared small targets typically consist of only 1-9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
-In this context, cross-task mechanisms that enable feature exchange between different tasks are more relevant to our work.
+In real-world scenarios, infrared small targets typically consist of only 1 to 9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
+In this context, cross-task mechanisms that enable semantic exchange between different tasks are more relevant to our work.
 These methods enhance MTL by facilitating feature exchange or distillation across different tasks, allowing each task to benefit from complementary signals of others.
-Methods like prediction-and-distillation network (PAD-Net) utilize attention-based mechanisms to distill multi-modal features effectively [31], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [32].
-While our work builds upon the foundational principles of multi-task learning and cross-task mechanisms, it distinguishes itself in the feature hard-exchange mechanism.
-Diverging from the soft fusion approach used in attention-based methods, we employ a feature hard-exchange mechanism for feature fusion.
-This simple yet effective strategy ensures a more robust integration of features from different tasks.
+Methods like prediction-and-distillation network utilize attention-based mechanisms to distill multi-modal features [34], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [35].
+While our research shares the underpinnings of multi-task learning and cross-task mechanisms, we pioneer a new approach to feature fusion.
+Instead of following the well-trodden path of soft fusion found in attention-based methods, our work introduces an innovative protocol–-a dynamic hard-exchange mechanism.
+To the best of our knowledge, this is the first work to explore the potential of dynamic hard-exchange of features between tasks.
 III The DenseSIRST Dataset
-Figure 2: The four images above show diverse backgrounds, including city, sea, cloud, and mountain. The five diagrams below illustrate various targets, including Point, Cluster, UAV, Plane, and Ship.
-Figure 3: A total of 12 dataset images are shown.
-In each set of images, the top figure shows the simulated dense small target, and the bottom figure shows the corresponding sky segmentation of the image.
+Figure 2: Illustration of selected images from our proposed DenseSIRST dataset.
+Each image pair consists of a simulated dense small target image (top) and its corresponding sky-segmented version (bottom).
+These images highlight the challenges posed by densely clustered targets within diverse backgrounds.
 III-A Motivation and Rationale
-The detection of small, low-contrast targets against complex backgrounds in infrared imagery is a significant challenge that impacts the performance of computer vision algorithms.
-While existing datasets have contributed to the advancement of this field, they often suffer from limitations such as sparse target distribution and a focus on larger target sizes, which do not adequately represent real-world scenarios.
-To address these limitations, we introduce the DenseSIRST dataset, a richly annotated collection of infrared images featuring densely populated small targets.
-DenseSIRST not only increases the number of small targets per image but also encompasses a variety of target sizes and complex clustering, crucial for training and evaluating algorithms under realistic and challenging conditions.
-TABLE I: Main characteristics of several popular SIRST datasets.
+The detection of small, low-contrast targets against complex backgrounds in infrared imagery poses a significant challenge.
+Existing datasets, though instrumental in progressing the field, exhibit critical deficiencies.
+As shown in Table I, well-regarded datasets such as SIRST V1 [15], SIRST V2 [3], IRSTD1K [10], and SIRSTAUG [36] predominantly feature sparsely distributed targets and lack per-pixel background annotations.
+This stands in stark contrast to the dense, diminutive targets frequently encountered in real-world scenarios.
+To bridge this gap, we present DenseSIRST: a meticulously annotated dataset of infrared images teeming with densely clustered, small targets, which is a more realistic and challenging benchmark for algorithm training and evaluation.
+Fig.
+2 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations.
+These annotations provide valuable contextual information for developing and evaluating detection algorithms.
+TABLE I: Comparison of the DenseSIRST dataset with other popular datasets used in infrared small target detection.
 Datasets Image Type Annotation Type Image Number Target Number Average Target Area Sparse or Clustered Background Semantic Annotation
-SIRST [15] Real Pixel 427 533 23 Sparse
+SIRST V1 [15] Real Pixel 427 533 23 Sparse
 SIRST V2 [3] Real Pixel + BBox + Point 1024 648 24 Sparse
@@
 DenseSIRST (Ours) Real + Synthetic Pixel + BBox + Point 1024 13655 6 Cluster
-Table I compares DenseSIRST with several well-regarded datasets, such as SIRST [15], SIRSTv2 [4], IRSTD1K [12], and SIRSTAUG [33].
-These datasets predominantly feature sparsely distributed targets and lack comprehensive annotations for clustered small targets.
-In contrast, DenseSIRST distinguishes itself through:
-  1. 1.
-    Hybrid Image Composition: Combining real and synthetic targets, DenseSIRST enhances both the realism and diversity of the training data.
-  2. 2.
-    Target Density and Diversity: With an average target area significantly smaller than existing datasets and a higher number of targets per image, DenseSIRST provides a challenging scenario for detection algorithms.
-  3. 3.
-    Varied Background Scenarios: The dataset includes images from diverse backgrounds such as urban, mountainous, maritime, and cloudy scenes, as shown in Fig.
-2, each contributing to the complexity of the detection tasks.
-Fig.
-3 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations.
-These annotations provide valuable contextual information for developing and evaluating detection algorithms.
+Figure 3: Statistical characteristics of our DenseSIRST dataset, underscoring the detection challenges it presents.
+(a) The distribution of target sizes, with the circle sizes in the scatter plot indicating the prevalence of each size category.
+(b) The distribution of local contrasts for small targets in the dataset, demonstrating that the dataset encompasses a broad range of contrasts, down to a minimum of 0.15.
+(c) The brightness distribution of small targets in the dataset, highlighting the wide spectrum from dark to light targets this dataset offers for detection tasks.
 III-B Target Synthesis and Composition
@@
 We curated this dataset from diverse real-world scenarios, supplemented with synthetic targets designed for physical plausibility and visual realism.
-The dataset leverages the existing SIRSTv2, creating a broad library of infrared small targets.
+The dataset leverages our SIRST V2 dataset, creating a broad library of infrared small targets.
 Each image is meticulously designed to emulate the densely populated surveillance scenarios.
@@
 This modification is governed by the following equation:
-$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1]$ (1)
-Here, $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location.
+$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1],$ (1)
+where $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location.
 Different from mixup, its weight is set to $1$ for enhancing the reality of the synthesized target.
@@
 $G(x,y)=e^{-\frac{{(x-\rho_{x}w)^{2}}}{{2\sigma_{x}^{2}}}-\frac{{(y-\rho_{y}h)^{2}}}{{2\sigma_{y}^{2}}}},$ (2)
-The relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$.
+where the relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$.
 These values, when multiplied by the target bounding box’s width $w$ and height $h$, yield the actual pixel offsets, ensuring appropriate scaling with the target size.
@@
 This randomness guarantees varying degrees of edge blurring among the simulated targets, reinforcing their realistic appearance.
-To further enhance variability, each target undergoes a rotational transformation. The Gaussian matrix’s rotation $\theta$ is applied as:
+To further enhance variability, each target undergoes a rotational transformation.
+The Gaussian matrix’s rotation $\theta$ is applied as:
 $\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x-\mu_{x}\\ y-\mu_{y}\end{bmatrix},$ (3)
-Here, $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
+where $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
 III-C Dataset Description and Characteristics
-III-C1 Composition of the Dataset
+III-C1 Dataset Composition
 The DenseSIRST dataset comprises 1024 infrared images, with a total of 13,655 densely clustered small targets.
 The images in the DenseSIRST dataset cover a wide range of realistic scenarios, including urban areas, mountainous regions, maritime environments, and cloudy scenes.
-This diversity enables the development and assessment of infrared small target detection algorithms that can generalize well to various real-world applications.
-The division of training set, validation set, and test set is aligned with the SIRSTv2 dataset.
+This diversity enables the development and assessment of detection algorithms that can generalize well to various real-world applications.
+The division of training, validation, and test set is aligned with the SIRST V2 dataset.
 III-C2 Annotations and Labels
@@
 Point annotations mark the center of each target, allowing for the evaluation of localization accuracy.
-III-C3 Statistical Characteristics of the Dataset
-The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection.
-This part offers a detailed analysis of the dataset’s key characteristics, including the distribution of image sizes, target sizes, local contrast, and brightness levels.
-Figure 4: Figure (a) is the image resolution characteristic, the larger the range of circles in the scatter plot, the greater the number of images at that resolution.
-Figure (b) is the target size characteristic, and the larger the range of circles in the scatter plot, the greater the number of small targets of that size.
-Figure (c) shows the percentage of different local contrasts for small objects in the dataset, with the lowest local contrast being 0.15.
-Figure (d) shows the percentage of brightness distribution for small targets in the dataset, and shows that the brightness distribution is wide from dark targets to light targets.
-Image Size Distribution: The dataset, as depicted in Fig.
-4 (a), includes images with resolutions ranging from $320\times 256$ to $1280\times 1024$ pixels, covering various typical infrared camera resolutions encountered in real-world scenarios.
-This diversity is crucial for developing detection algorithms that can generalize across different resolutions.
+III-C3 Statistical Characteristics
+The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection. This part offers an analysis of the dataset’s key characteristics.
 Target Size Distribution: Most targets, as indicated in Fig.
@@
 The $3\times 3$ pixels size is the most frequent, aligning with the characteristics of small target detection in infrared imagery.
-With a total of 13,655 annotated targets, the dataset provides a rich set of examples for training and evaluation.
-Local Contrast Distribution: The vast majority (90%) of targets have local contrast values below $2$, as shown in Fig.
+Local Contrast Distribution: The 90% targets have local contrast below $2$, as shown in Fig.
 3 (c).
@@
 3 (d) reveals that only a minority of targets are the brightest points in their respective images.
-This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection in infrared images.
+This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection.
 In conclusion, the DenseSIRST dataset, with its wide range of image dimensions, prevalence of small target sizes, and low local contrast, provides a realistic and challenging benchmark for infrared small target detection algorithms.
@@
 After determining the appropriate pasting position, the Gaussian technique is used to integrate the small targets into the new background.
-The proposed BAFE-Net adopts a fully convolutional one-stage object detector (FCOS) [34] as its host network.
-FCOS is a simple yet effective anchor-free object detection framework that has demonstrated superior performance in various computer vision tasks.
-By leveraging the inherent advantages of FCOS, such as its ability to handle objects of different scales and aspect ratios without the need for predefined anchor boxes, BAFE-Net aims to address the challenges of detecting small, densely clustered targets in infrared imagery.
-The overall architecture of BAFE-Net is illustrated in Fig.
-5.
-In addition to the original detection head present in the baseline FCOS, BAFE-Net introduces a segmentation head within the BAFE-Head module.
+The overall architecture of BAFE-Net, which adopts the fully convolutional one-stage object detector (FCOS) [37] as its host network, is illustrated in Fig.
+4.
+In addition to the original detection head present in the baseline, BAFE-Net introduces a segmentation head within the BAFE-Head module.
 The primary purpose of incorporating this segmentation head is to enable explicit modeling of background semantics, which is crucial for enhancing the network’s ability to distinguish between real targets and background clutter in infrared scenes.
-By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information, ultimately leading to improved detection performance.
-In addition to the architecture modifications, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP).
-This technique is used during the training phase to augment the training dataset with realistically integrated synthetic targets.
-BAG-CP strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, which significantly enhances the training process.
+By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information.
+Furthermore, BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to facilitate the bidirectional flow of information between the detection and segmentation branches.
+This dynamic exchange allows the network to adaptively leverage the most relevant features from each task, leading to improved detection performance.
+To address the challenge of limited target samples, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP).
+This technique strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, ensuring that the pasted targets are only placed in sky regions.
 The use of Gaussian smoothing at the edges of the pasted targets ensures seamless integration, avoiding artificial boundary artifacts that could mislead the learning algorithm.
-This context-aware synthesis and integration approach allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds.
-By training the network to recognize and handle these challenging conditions effectively, BAFE-Net sets a new standard for performance in infrared small target detection.
-IV-B Cross-Task Feature Exchange Module
-During the feature extraction process, partial feature exchange is performed between segmentation and classification tasks within this network.
-The segmentation and classification tasks have different focuses when learning features: segmentation emphasizes local details and contextual information, whereas classification emphasizes global features.
-Despite their differing objectives, there is a potential synergistic relationship between segmentation and classification tasks.
-By exchanging features, information can be shared between the two tasks, promoting collaborative learning and enhancing the feature representation capabilities of each task.
-Segmentation can assist classification in better understanding the structure and semantics of the image, while classification can help segmentation in distinguishing boundaries and details of different categories.
-Specifically, segmentation features can leverage global information from classification features to improve overall understanding, while classification features can utilize local detail information from segmentation features to enhance precision.
-Mutual learning between segmentation and classification features can reduce the risk of overfitting and improve the model’s performance in practical applications.
-Feature exchange enables the model to perform more robustly when facing different tasks.
-It allows the model to better adapt to various changes and uncertainties, thereby enhancing its generalization ability across different scenarios.
-Feature exchange has been proven feasible for several reasons: 1) Complementarity of features: Although segmentation and classification features are different, they are extracted from the same input image and thus possess a degree of complementarity.
-By exchanging features, each task can utilize useful information from the other task, thereby improving overall performance.
-2) Compatibility of features: Deep learning models use continuous and differentiable operations, such as convolution and pooling, when extracting features.
-Through feature exchange, these operations can still ensure the coherence and differentiability of features, thus not disrupting the training process of the model.
-3) Support from existing research: Some studies have already shown that in multi-task learning, sharing or exchanging features can improve the overall performance of tasks.
-For example, in the Changer [35] paper, channel and spatial exchanges were used and proven feasible.
-Feature exchange primarily consists of two methods: channel exchange and spatial exchange.
-Channel exchange involves directly swapping features at the channel level between different feature maps, while spatial exchange is achieved by swapping features along the spatial dimensions of the feature maps.
-In the Cross-Task Feature Exchange Module, channel exchange is employed, and experiments have shown that channel exchange performs better than spatial exchange.
-The experimental results will be presented in Section V.
-Assume ${x}_{1}$ and ${x}_{2}$ are the input tensors, each with a shape of $(N,C,H,W)$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width, respectively.
-Let $p$ be a parameter controlling the proportion of channel exchange.
-Define an exchange mask $E$ with a shape of $C$,
-${E}[i]=\begin{cases}1&\text{, if }i\mod\left(\frac{1}{p}\right)=0\\ 0\end{cases}$ (4)
-where $i$ is the channel index. Expand the exchange mask $E$ to the shape of $(N,C,H,W)$,
-${E}_{\text{expanded}}={E}\otimes{1}_{N\times H\times W}$ (5)
-where $\otimes$ denotes the expansion operation, and ${1}_{N\times H\times W}$ is a tensor of ones with a shape of $(N,H,W)$. Define the output tensors ${y}_{1}$ and ${y}_{2}$ ,
-${y}_{1}={x}_{1}\odot(1-{E}_{\text{expanded}})+{x}_{2}\odot{E}_{\text{expanded}}$ (6)
-${y}_{2}={x}_{2}\odot(1-{E}_{\text{expanded}})+{x}_{1}\odot{E}_{\text{expanded}}$ (7)
-where $\odot$ denotes element-wise multiplication.
-However, the features to be exchanged are cross-modal.
-Therefore, an Adapter is introduced to process these features before the exchange.
-Here, an MLP (Multi-Layer Perceptron) is used as the Adapter.
-The features to be exchanged are fed into the Adapter separately, and after processing, the features are exchanged.
-This approach addresses the cross-modal issue between the features.
+By augmenting the training dataset with realistically integrated synthetic targets, BAG-CP allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds.
+This context-aware synthesis and integration approach enables BAFE-Net to effectively recognize and handle these challenging conditions, ultimately setting a new standard for performance in infrared small target detection.
+IV-B Dynamic Cross-Task Feature Exchange Head
+For infrared small target detection, the scarcity of intrinsic target features and the abundance of similar background distractors pose significant challenges.
+To address this, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation.
+The motivation behind this cross-task feature exchange is twofold:
+  1. 1.
+    Complementary Nature of Tasks: Although target detection and background segmentation have different objectives, they share a common goal of understanding the scene.
+By exchanging features between these tasks, each task can benefit from the complementary information learned by the other, leading to a more comprehensive understanding of the image.
+  2. 2.
+    Synergistic Relationship: Target detection focuses on identifying the presence and location of small targets, while background segmentation emphasizes the contextual information and local details of the scene.
+By sharing features between these tasks, target detection can leverage the semantic understanding of the background to better distinguish targets from distractors, while background segmentation can utilize the localized target information to refine its segmentation boundaries.
+IV-B1 Dynamic Channel Selection
+To further optimize cross-task feature exchange, we propose a dynamic selection mechanism that determines the most relevant feature channels for exchange based on their importance to task performance.
+The importance of each channel is estimated using an attention mechanism, which assigns weights to channels according to their contribution.
+Given an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature map, respectively, we first compute a channel-wise attention weight vector $\mathbf{m}\in\mathbb{R}^{C}$ by a mechanism inspired by the design of our previous work Large Selective Kernel Network (LSKNet) [38].
+This is achieved by applying a series of operations to $\mathbf{X}$, including local feature extraction, channel-wise statistics computation, and weighted fusion.
+The process can be summarized as follows:
+$\mathbf{m}$ $=\sigma\left(\mathbf{X}\odot\left(\mathcal{F}(\mathbf{X})\odot\sigma\left(\mathcal{S}\left(\text{avg}(\mathbf{X})\oplus\max(\mathbf{X})\right)\right)\right)\right),$ (4)
+where $\odot$ denotes element-wise multiplication, $\oplus$ denotes the concatenation operation, and $\sigma$ is the Sigmoid function.
+$\max$ and avg are the maximum and average pooled spatial feature descriptors.
+$\mathcal{S}$ is a convolution operation that reduces the dimensionality of concatenated features.
+$\mathcal{F}$ is also a convolution operation applied to $\mathbf{X}$ , extracting local features.$\mathbf{m}\in\mathbb{R}^{C}$ is a one-dimensional tensor, and $C$ is the number of channels.
+Each element $\mathbf{m}$ represents the attention logits for the $i$-th channel.
+For each sample, the top $K$ channels with the highest weights are selected, where $K$ is determined as a fraction $p$ of the total number of channels $C$:
+$K$ $=C\times p,$ (5)
+${V}_{\text{topk}},{I}_{\text{topk}}$ $=\mathrm{TopK}({\mathbf{m}},K),$
+where ${V}_{\text{topk}}\in\mathbb{R}^{K}$ represents the top $K$ largest weight values for each sample.
+${I}_{\text{topk}}\in\mathbb{R}^{K}$ represents the corresponding channel indices for these weights.
+This process dynamically selects the channels for exchange, improving the effectiveness of cross-task feature sharing.
+IV-B2 Cross-Task Feature Hard-Exchange
+Within BAFE-Net, we introduce a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks.
+Feature hard-exchange allows for the direct transfer of learned features from one task to another.
+By swapping entire feature channels, the receiving task has access to the explicit representation learned by the other task, facilitating a more direct integration of complementary information.
+To perform feature exchange, let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ be the input tensors from the target detection and background segmentation tasks, respectively, each with a shape of $(C,H,W)$, where $C$ is the number of channels, and $H$ and $W$ are the height and width.
+We define an exchange mask $\mathbf{e}\in\mathbb{R}^{C}$:
+$\mathbf{e}[i]=\begin{cases}1,&\mathrm{if}\ i\in I_{\text{topk}}\\ 0,&\mathrm{otherwise}\end{cases},$ (6)
+where $i$ represents the channel index. The exchange mask $\mathbf{e}$ is then expanded to the shape of $(C,H,W)$:
+$\mathbf{E}_{d}=\mathbf{e}\otimes\mathbf{1}_{H\times W},$ (7)
+where $\otimes$ denotes the expansion operation, and $\mathbf{1}_{H\times W}$ is a tensor of ones with a shape of $(H,W)$.
+$\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ can obtain $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ respectively through Eq.
+(4) to Eq.
+(7).
+The output tensors $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ after the hard feature exchange are computed as:
+$\mathbf{Y}_{1}=\mathbf{X}_{1}\odot(\mathbf{1}-\mathbf{E}_{d1})+\mathcal{A}\left(\mathbf{X}_{2}\odot\mathbf{E}_{d2}\right),$ (8)
+$\mathbf{Y}_{2}=\mathbf{X}_{2}\odot(\mathbf{1}-\mathbf{E}_{d2})+\mathcal{A}\left(\mathbf{X}_{1}\odot\mathbf{E}_{d1}\right),$ (9)
+where $\odot$ denotes element-wise multiplication, $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ represent the output features, and $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ are the dynamic masks.
+To address the cross-modal nature of features, we introduce an adapter module $\mathcal{A}$, implemented as a multi-layer perceptron (MLP), which projects features into a common space before exchange.
+This design enables effective cross-modal interaction while preserving modality-specific characteristics.
 IV-C Background-Aware Gaussian Copy-Paste
@@
 As illustrated in Fig.
-6, The BAG-CP process unfolds as follows:
+5, The BAG-CP process unfolds as follows:
 Prerequisite: Semantic Annotation of Background: Before performing the copy-paste operation, it is crucial to have the image segmented into different semantic regions.
@@
   1. 1.
-    Ablation Study on Background-Aware Gaussian Copy-Paste. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
+    Ablation Study on BAG-CP. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
   2. 2.
-    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlights the contributions of each component in BAFE-Net.
+    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlight the contributions of each component in BAFE-Net.
   3. 3.
@@
 1) Dataset: During the experimental phase of our research, we conducted tests on the proposed dataset DenseSIRST.
-2) Implementation Details: Our network was trained on an NVIDIA RTX 3090 GPU platform with 24GB of memory.
-In our experiments, we employed the DeepInfrared toolkit for object detection.
-Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion.
+2) Implementation Details: Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion.
 We choose Focal Loss as the loss function of the classification branch, IoU Loss as the loss function of the regression branch, and Cross Entropy Loss as the loss function of the segmentation branch.
-We choose DAdaptAdam as the optimizer, lr is set to 1.0, and weight decay is set to 0.05.
+We choose DAdaptAdam as the optimizer, learning rate is set to 1.0, and weight decay is set to 0.05.
 Considering training and convergence time, we trained for 20 epochs within the DeepInfrared framework.
-Through experimentation, we observed that different parameter settings affect the detection efficacy of dense small objects.
-Thus, while ensuring basic fairness, we adjusted the parameters of different methods to obtain optimal detection results.
-3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment. They are defined as:
-$\text{mAP}=\frac{1}{N}\sum_{i=1}^{N}\text{AP}(i),$ (8)
-[…]
-$\text{Recall}=\frac{TP}{TP+FN},$ (10)
-Here, $N$ is the total number of samples, $TP$, $FP$, and $FN$ denote the number of true positive, false positive, and false negative instances, respectively.
-Precision $p(r)$ is defined as the proportion of true positive detections out of the total number of positive detections at a given recall level $r$.
-TABLE II: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net
-Strategy Module DenseSIRST
-Copy-Paste Cluster Gaussian Background-Aware mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) 0.225 0.311 0.180 0.314
-(b) $\checkmark$ 0.236 0.322 0.203 0.303
-(c) $\checkmark$ $\checkmark$ 0.242 0.329 0.205 0.335
-(d) $\checkmark$ $\checkmark$ 0.245 0.321 0.211 0.338
-(e) $\checkmark$ $\checkmark$ $\checkmark$ 0.250 0.319 0.215 0.315
-(f) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.258 0.340 0.220 0.340
-TABLE III: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
-Strategy Module DenseSIRST
-Seg Head Feature Exchange Adapter mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) 0.225 0.311 0.180 0.314
-(b) $\checkmark$ 0.246 0.324 0.183 0.304
-(c) $\checkmark$ $\checkmark$ 0.254 0.321 0.199 0.310
-(d) $\checkmark$ $\checkmark$ $\checkmark$ 0.262 0.318 0.208 0.305
-TABLE IV: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
-Strategy Hidden layers DenseSIRST
-mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) 8 0.222 0.305 0.206 0.308
-(b) 16 0.243 0.318 0.214 0.310
-(c) 32 0.256 0.329 0.218 0.312
-(d) 64 0.266 0.336 0.230 0.323
-TABLE V: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
-Strategy Exchange Method DenseSIRST
-mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) Channel 0.266 0.336 0.230 0.323
-(b) Spatial 0.223 0.304 0.218 0.322
-(c) Channel + Spatial 0.246 0.310 0.199 0.314
-(d) Spatial + Channel 0.224 0.310 0.201 0.328
-TABLE VI: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
-Strategy ResNet DenseSIRST
-mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) ResNet18 0.258 0.328 0.233 0.341
-(b) ResNet34 0.258 0.330 0.226 0.332
-(c) ResNet50 0.270 0.332 0.236 0.329
-(d) ResNet101 0.256 0.324 0.241 0.336
-TABLE VII: Comparison with Other State-of-the-art methods on DenseSIRST.
+3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment.
+TABLE II: Comparison with Other State-of-the-art methods on DenseSIRST.
 Method Backbone DenseSIRST
-mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ Flops $\downarrow$ Params $\downarrow$
+FLOPs $\downarrow$ Params $\downarrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
               One-stage
-FCOS [34] ResNet50 0.232 0.315 0.204 0.324 50.291G 32.113M
-SSD [36] 0.211 0.421 0.178 0.424 87.552G 23.746M
-GFL [37] ResNet50 0.253 0.332 0.230 0.317 52.296G 32.258M
-ATSS [38] ResNet50 0.248 0.327 0.202 0.326 51.504G 32.113M
-CenterNet [39] ResNet50 0.000 0.000 0.000 0.000 50.278G 32.111M
-PAA [40] ResNet50 0.255 0.545 0.228 0.551 51.504G 32.113M
-PVT-T [41] 0.109 0.481 0.093 0.501 41.623G 21.325M
-RetinaNet [42] ResNet50 0.114 0.510 0.086 0.523 52.203G 36.330M
-EfficientDet [43] 0.099 0.433 0.072 0.419 34.686G 18.320M
-TOOD [13] ResNet50 0.256 0.355 0.226 0.342 50.456G 32.018M
-VFNet [44] ResNet50 0.253 0.336 0.214 0.336 48.317G 32.709M
-YOLOF [45] ResNet50 0.091 0.009 0.002 0.009 25.076G 42.339M
-AutoAssign [46] ResNet50 0.255 0.354 0.180 0.314 50.555G 36.244M
-DyHead [47] ResNet50 0.249 0.335 0.189 0.328 27.866G 38.890M
+SSD [39] 87.552G 23.746M 0.211 0.421 0.178 0.424
+RetinaNet [40] ResNet50 52.203G 36.330M 0.114 0.510 0.086 0.523
+YOLOv3 [41] 50.002G 61.949M 0.233 0.424 0.207 0.413
+CenterNet [42] ResNet50 50.278G 32.111M 0.138 0.316 0.124 0.317
+FCOS [37] ResNet50 50.291G 32.113M 0.232 0.315 0.204 0.324
+ATSS [43] ResNet50 51.504G 32.113M 0.248 0.327 0.202 0.326
+CentripetalNet [44] 0.491T 0.206G 0.244 0.259 0.201 0.244
+AutoAssign [45] ResNet50 50.555G 36.244M 0.255 0.354 0.180 0.314
+GFL [46] ResNet50 52.296G 32.258M 0.264 0.367 0.230 0.317
+PAA [47] ResNet50 51.504G 32.113M 0.255 0.545 0.228 0.551
+VFNet [48] ResNet50 48.317G 32.709M 0.253 0.336 0.214 0.336
+PVT-T [49] 41.623G 21.325M 0.109 0.481 0.093 0.501
+YOLOF [50] ResNet50 25.076G 42.339M 0.091 0.009 0.002 0.009
+YOLOX [51] 8.578G 8.968M 0.210 0.341 0.180 0.331
+TOOD [12] ResNet50 50.456G 32.018M 0.256 0.355 0.226 0.342
+DyHead [52] ResNet50 27.866G 38.890M 0.249 0.335 0.189 0.328
+DDOD [53] ResNet50 46.514G 32.378M 0.253 0.335 0.230 0.351
+RTMDet [54] 51.278G 52.316M 0.229 0.349 0.212 0.350
+EfficientDet [55] 34.686G 18.320M 0.146 0.464 0.094 0.517
 Two-stage
-Faster R-CNN [48] ResNet50 0.091 0.022 0.015 0.029 0.759T 33.035M
-Cascade R-CNN [49] ResNet50 0.136 0.188 0.139 0.194 90.978G 69.152M
-Dynamic R-CNN [50] ResNet50 0.184 0.235 0.111 0.190 63.179G 41.348M
-Grid R-CNN [51] ResNet50 0.091 0.018 0.025 0.037 0.177T 64.467M
-Libra R-CNN [52] ResNet50 0.141 0.142 0.085 0.120 63.990G 41.611M
+Faster R-CNN [56] ResNet50 0.759T 33.035M 0.091 0.022 0.015 0.029
+Cascade R-CNN [57] ResNet50 90.978G 69.152M 0.136 0.188 0.139 0.194
+Grid R-CNN [58] ResNet50 0.177T 64.467M 0.156 0.122 0.104 0.190
+Libra R-CNN [59] ResNet50 63.990G 41.611M 0.141 0.142 0.085 0.120
+TridentNet [60] ResNet50 0.759T 33.035M 0.091 0.009 0.014 0.021
+SABL [61] ResNet50 0.125T 42.213M 0.124 0.104 0.104 0.171
+Dynamic R-CNN [62] ResNet50 63.179G 41.348M 0.184 0.235 0.111 0.190
 End2End
-DETR [53] ResNet50 0.000 0.000 0.000 0.000 24.940G 41.555M
-Deformable DETR [54] ResNet50 0.024 0.016 0.018 0.197 51.772G 40.099M
-DAB-DETR [55] ResNet50 0.005 0.054 0.000 0.001 28.939G 43.702M
-Conditional DETR [56] ResNet50 0.000 0.000 0.000 0.001 27.143G 40.297M
-Sparse R-CNN [57] ResNet50 0.183 0.572 0.154 0.614 45.274G 0.106G
-$\star$ BAFE-Net (Ours) ResNet50 0.270 0.332 0.236 0.329 69.114G 35.329M
-V-B Ablation Study
-V-B1 Ablation Study on Background-Aware Gaussian Copy-Paste
-In this study, we initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting.
-Our aim is to empirically validate the enhancements to detection performance brought about by these steps.
-To achieve this, we explore various combinations of the aforementioned components and conduct experimental validation of the results.
-The experimental outcomes, as outlined in Table II, generate several key insights, providing a holistic understanding of the impact of each module.
-Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study.
-The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
-Further refinement through clustering of pasted objects also yields substantial improvements.
-This is evident when comparing strategies (b) to (c), and (d) to (e), respectively, indicating the considerable value of object clustering.
-Applying Gaussian blur to the edges of the pasted objects, as a means to enhance their realistic appearance, also proves to be significantly beneficial in improving detection performance.
-This is substantiated when comparing strategies (b) with (d), and (c) with (e).
-Finally, the culmination of all modules in a Background-Aware Gaussian Copy-Paste context, as depicted by strategy (f), results in further enhancements in detection capability.
-This is demonstrated with an additional increase of 0.8% on the DenseSIRST dataset when comparing strategy (e) with strategy (f).
-V-B2 Ablation Study on the Scheme of Background-Aware Feature Exchange
-The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table III.
-To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
-Table III encapsulates the results of our experimental study and yields several crucial insights.
-Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research.
-Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head.
-This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
-Further, juxtaposition of Strategy (b) and Strategy (c) highlights the contribution of the Feature Exchange in the interaction mode between the detection head and the segmentation head.
-This interaction led to a notable improvement in detection performance by 1.1% on the DenseSIRST dataset.
-Upon comparing Strategy (c) and Strategy (d), it becomes evident that introducing feature adapation prior to the Feature Exchange further amplifies detection performance.
-As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration.
-Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity.
-To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  IV.
-The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance.
-However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially.
-Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
-V-B3 Ablation Study on Feature Exchange Mechanism
-In this part, a rigorous exploration of various Feature Exchange Mechanisms was undertaken.
-This encompassed diverse mechanisms such as Channel Exchange and Spatial Exchange, along with a multitude of their combinations.
-The quest was to ascertain the most efficacious method for Feature Exchange.
-Channel Exchange, as a mechanism, facilitates the direct interplay at the channel level amongst different features.
-In contrast, Spatial Exchange is implemented by orchestrating exchanges along the spatial dimensions inherent in the feature maps.
-The exchange ratio was judiciously set at 0.5 to maintain a balance.
-As delineated in Table V, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested.
-The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
-V-B4 Impact of Backbone Depth:
-For infrared small target detection, the depth of the backbone network emerges as a critical factor.
-A shallower network, while computationally efficient, might extract features with insufficient semantic richness, thereby impinging on the overall detection performance.
-Conversely, an excessively deep network might overwhelm the delicate features of smaller targets amidst the background, consequently leading to a performance degradation.
-We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net.
-The results of these experiments are summarized in Table VI.
-Notably, ResNet50 emerged as the most suitable backbone for the BAFE-Net.
-V-C Comparison with State-of-the-Arts
+DETR [63] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000
+Sparse R-CNN [64] ResNet50 45.274G 0.106G 0.183 0.572 0.154 0.614
+Deformable DETR [65] ResNet50 51.772G 40.099M 0.024 0.016 0.018 0.197
+Conditional DETR [66] ResNet50 27.143G 40.297M 0.000 0.000 0.000 0.001
+DAB-DETR [67] ResNet50 28.939G 43.702M 0.005 0.054 0.000 0.001
+$\star$ BAFE-Net (Ours) ResNet18 57.654G 22.31M 0.283 0.335 0.233 0.325
+$\star$ BAFE-Net (Ours) ResNet50 71.639G 35.626M 0.274 0.342 0.248 0.338
+Figure 6: Feature visualization comparison. Compared with other methods, the proposed BAFE-Net shows superior performance with higher detection of small objects and fewer false alarms.
+V-B Comparison with State-of-the-Arts
 We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall.
-The experimental results are presented in Table VII.
+The experimental results are presented in Table II.
 End-to-end (End2End) methods, while theoretically appealing due to their NMS-free style, demonstrated suboptimal performance.
@@
 The superior performance of BAFE-Net validates the efficacy of our approach in accurately detecting small infrared targets amidst dense conditions.
-V-D Visual Analysis
+TABLE III: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net
+Strategy Module DenseSIRST
+Copy-Paste Cluster Gaussian Background-Aware mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) 0.225 0.311 0.180 0.314
+(b) $\checkmark$ 0.236 0.322 0.203 0.303
+(c) $\checkmark$ $\checkmark$ 0.242 0.329 0.205 0.335
+(d) $\checkmark$ $\checkmark$ 0.245 0.321 0.211 0.338
+(e) $\checkmark$ $\checkmark$ $\checkmark$ 0.250 0.319 0.215 0.315
+(f) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.258 0.340 0.220 0.340
+TABLE IV: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
+Strategy Module DenseSIRST
+Seg Head Feature Exchange Adapter mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) 0.225 0.311 0.180 0.314
+(b) $\checkmark$ 0.246 0.324 0.183 0.304
+(c) $\checkmark$ $\checkmark$ 0.254 0.321 0.199 0.310
+(d) $\checkmark$ $\checkmark$ $\checkmark$ 0.262 0.318 0.208 0.305
+TABLE V: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
+Strategy Hidden layers DenseSIRST
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) 8 0.222 0.305 0.206 0.308
+(b) 16 0.243 0.318 0.214 0.310
+(c) 32 0.256 0.329 0.218 0.312
+(d) 64 0.266 0.336 0.230 0.323
+TABLE VI: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
+Strategy Exchange Method DenseSIRST
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) Channel 0.266 0.336 0.230 0.323
+(b) Spatial 0.223 0.304 0.218 0.322
+(c) Channel + Spatial 0.246 0.310 0.199 0.314
+(d) Spatial + Channel 0.224 0.310 0.201 0.328
+TABLE VII: Ablation study on the impact of feature selection on detection performance in Channel-wise Feature Exchange
+Strategy Feature Selection DenseSIRST
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) Random Selection 0.255 0.322 0.172 0.317
+(b) Fixed Selection 0.273 0.333 0.230 0.323
+(c) Dynamic Selection 0.283 0.335 0.233 0.325
+TABLE VIII: Ablation study on the impact of the number of exchange channels on detection performance in Top-K selection
+Strategy P DenseSIRST
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) 1/2 0.283 0.335 0.233 0.325
+(b) 1/4 0.241 0.316 0.230 0.331
+(c) 3/4 0.242 0.325 0.204 0.310
+(d) 1/8 0.238 0.312 0.218 0.323
+(e) 3/8 0.255 0.327 0.203 0.331
+(f) 5/8 0.259 0.325 0.201 0.328
+(g) 7/8 0.238 0.316 0.214 0.329
+(h) 1/16 0.245 0.326 0.213 0.301
+(i) 1/32 0.236 0.319 0.212 0.312
+TABLE IX: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
+Strategy ResNet DenseSIRST
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) ResNet18 0.283 0.335 0.233 0.341
+(b) ResNet34 0.281 0.354 0.226 0.332
+(c) ResNet50 0.274 0.342 0.248 0.338
+(d) ResNet101 0.261 0.317 0.241 0.336
+V-C Visual Analysis
 To further validate the effectiveness of our proposed methods, we conduct a visual analysis by comparing the feature activation maps of our approach with those of the top-performing competing methods in the quantitative evaluation.
@@
 As evident from Fig.
-7, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference.
+6, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference.
 This can be attributed to the background-aware mechanism employed by BAFE-Net, which learns to differentiate between target and background regions during the feature exchange process.
 By minimizing the activation levels in the background areas, our method demonstrates enhanced robustness against false positives, leading to a higher detection precision.
-Figure 7: Feature visualization comparison. Compared with other methods, the proposed BAFE-Net shows superior performance with higher detection of small objects and fewer false alarms.
+V-D Ablation Study
+V-D1 Ablation Study on Background-Aware Gaussian Copy-Paste
+We initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting.
+The experimental outcomes, as outlined in Table III, generate several key insights, providing a holistic understanding of the impact of each module.
+Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study.
+The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
+Further refinement through clustering of pasted objects also yields substantial improvements.
+This is evident when comparing strategies (b) to (c), and (d) to (e), respectively, indicating the considerable value of object clustering.
+Applying Gaussian blur to the edges of the pasted objects, as a means to enhance their realistic appearance, also proves to be significantly beneficial in improving detection performance.
+This is substantiated when comparing strategies (b) with (d), and (c) with (e).
+Finally, the culmination of all modules in a Background-Aware Gaussian Copy-Paste context, as depicted by strategy (f), results in further enhancements in detection capability.
+This is demonstrated with an additional increase of 0.8% on the DenseSIRST dataset when comparing strategy (e) with strategy (f).
+V-D2 Ablation Study on the Scheme of Background-Aware Feature Exchange
+The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table IV.
+To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
+Table IV encapsulates the results of our experimental study and yields several crucial insights.
+Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research.
+Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head.
+This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
+Further, juxtaposition of Strategy (b) and Strategy (c) highlights the contribution of the Feature Exchange in the interaction mode between the detection head and the segmentation head.
+This interaction led to a notable improvement in detection performance by 1.1% on the DenseSIRST dataset.
+Upon comparing Strategy (c) and Strategy (d), it becomes evident that introducing feature adapation prior to the Feature Exchange further amplifies detection performance.
+As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration.
+Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity.
+To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  V.
+The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance.
+However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially.
+Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
+V-D3 Ablation Study on Feature Exchange Mechanism
+In this part, a rigorous exploration of various Feature Exchange Mechanisms was undertaken.
+This encompassed diverse mechanisms such as Channel Exchange and Spatial Exchange, along with a multitude of their combinations.
+The quest was to ascertain the most efficacious method for Feature Exchange.
+Channel Exchange, as a mechanism, facilitates the direct interplay at the channel level amongst different features.
+In contrast, Spatial Exchange is implemented by orchestrating exchanges along the spatial dimensions inherent in the feature maps.
+The exchange ratio was judiciously set at 0.5 to maintain a balance.
+As delineated in Table VI, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested.
+The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
+V-D4 Ablation Study in Dynamic Channel Selection
+In this section, we first conduct a rigorous exploration of feature selection methods.
+This includes random selection, fixed selection, and dynamic selection.
+The task is to determine the most effective feature selection method.
+1) Randomly Selection: This method selects feature channels randomly, which may lead to suboptimal feature exchanges.
+In some cases, it could even select features that hinder the task, ultimately reducing model performance.
+2) Fixed Selection: While fixed selection provides consistency during training and eliminates uncertainty, it lacks the adaptability to dynamic input features, thus limiting the model’s performance in complex or varied scenarios.
+3) Dynamic Selection: dynamic selection maximizes the performance of the model and reduces unnecessary information redundancy by selecting the most useful feature channels for exchange.
+The exchange ratio is judiciously set to 0.5 to maintain a balance.
+As shown in Table VII, the advantages of dynamic selection are obvious.
+Empirical evidence highlights the robustness and optimality of dynamic selection, thus confirming its adoption as the preferred method for feature selection in our model.
+In addition, we conduct a rigorous exploration of the value of $p$ in Top-K selection. As shown in Table VIII, $p$ is optimal when it is 1/2.
+V-D5 Impact of Backbone Depth
+For infrared small target detection, the depth of the backbone network emerges as a critical factor.
+A shallower network, while computationally efficient, might extract features with insufficient semantic richness, thereby impinging on the overall detection performance.
+Conversely, an excessively deep network might overwhelm the delicate features of smaller targets amidst the background, consequently leading to a performance degradation.
+We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net.
+The results of these experiments are summarized in Table IX.
+Notably, ResNet18 has the highest $mAP_{07}$ metric and ResNet50 has the highest $mAP_{12}$ metric.
 VI Conclusion

### v3

@@
-Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset
+Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection
 Abstract
-Infrared small target detection poses unique challenges due to the scarcity of intrinsic target features and the abundance of similar background distractors.
-We argue that background semantics play a pivotal role in distinguishing visually similar objects for this task.
-To address this, we introduce a new task—-clustered infrared small target detection, and present DenseSIRST, a novel benchmark dataset that provides per-pixel semantic annotations for background regions, enabling the transition from sparse to dense target detection.
-Leveraging this dataset, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation.
-BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks.
-Furthermore, we propose the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which selectively pastes small targets into sky regions during training, avoiding the creation of false alarm targets in complex non-sky backgrounds.
-Extensive experiments validate the effectiveness of BAG-CP and BAFE-Net in improving target detection accuracy while reducing false alarms.
-The DenseSIRST dataset, code, and trained models are available at https://github.com/GrokCV/BAFE-Net.
-Index Terms:
-Infrared Small Target Detection; Clustered Target Detection; Dataset; Feature Exchange; Semantic Segmentation
+Infrared small target detection presents significant challenges due to the limited intrinsic features of the target and the overwhelming presence of visually similar background distractors.
+We contend that background semantics are critical for distinguishing between objects that appear visually similar in this context.
+To address this challenge, we propose a task—clustered infrared small target detection—and introduce DenseSIRST, a benchmark dataset that provides per-pixel semantic annotations for background regions.
+This dataset facilitates the shift from sparse to dense target detection.
+Building on this resource, we propose the Background-Aware Feature Exchange Network (BAFE-Net), a multi-task architecture that jointly tackles target detection and background semantic segmentation.
+BAFE-Net incorporates a dynamic cross-task feature hard-exchange mechanism, enabling the effective exchange of target and background semantics between the two tasks.
+Comprehensive experiments demonstrate that BAFE-Net significantly enhances target detection accuracy while mitigating false alarms.
+The DenseSIRST dataset, along with the code and trained models, is publicly available at https://github.com/GrokCV/BAFE-Net.
 I Introduction
 Figure 1: Illustration of the background semantics’ crucial role in infrared small target detection.
-The small target in the orange box represents a genuine target of interest against the sky background, while the small target in the blue box is a false alarm in the context of the building background.
-When considering only local regions, it is challenging to differentiate between clustered small targets and false alarms due to their similar appearances.
-By incorporating global contextual information, the distinction between real targets and false alarms becomes more apparent.
-Infrared imaging has emerged as a crucial tool in various domains, particularly in scenarios where long-distance imaging is required in the absence of visible light, such as night-time operations or environments with limited visibility [1].
-By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring, maritime target search [2], and rescue missions [3].
-However, in these applications, the targets of interest are often located at great distances, resulting in their appearance as extremely small-sized objects in the captured images.
-Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images.
-I-A Prior Works on Clustered Infrared Small Target Detection
-While the potential of infrared imaging has been recognized across various domains, existing infrared small target detection techniques are predominantly tailored for the detection of sparse targets [4, 5], which proves to be inadequate when faced with complex application scenarios such as drone swarm defense, multi-target tracking systems, and extensive maritime search and rescue operations.
-Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking clustered infrared small targets.
-As the distribution of small targets transitions from sparse to dense scenarios, existing infrared small target detection techniques face a multitude of challenges that drastically hinder their effectiveness.
-Firstly, methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies.
-Moreover, the reliance on local contrast features [6], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects.
-Furthermore, even state-of-the-art deep learning methods [7, 8, 9, 10] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, significantly adding to the complexity of detection tasks.
-For instance, Dai et al.
-[11] introduced a detection transformer designed for arbitrarily oriented objects, yet its applicability to dense infrared small target settings remains limited due to the lack of intrinsic target features.
-Therefore, current methods are plagued by inefficiencies when operating in dense settings, frequently resulting in missed detections and high false alarm rates.
-This calls for new solutions tailored to overcome the intricacies of dense infrared small target detection.
+When considering only local regions (as in the top-left image), it is challenging to differentiate between clustered small targets and false alarms due to their similar appearances.
+By incorporating global contextual information (as in the surrounding images), the distinction between real targets and false alarms becomes more apparent.
+From the global images, it can be seen that the sky background is highlighted in red.
+The small target in the blue box represents a genuine target of interest against the sky background, while the small target in the red box is a false alarm in the context of the building background.
+Infrared small target detection has emerged as a fundamental challenge in computer vision, with critical applications spanning military surveillance, drone detection, search and rescue missions, and night operations.
+The task involves identifying and localizing diminutive objects that appear as few pixels in infrared imagery, often captured under adverse weather conditions or limited visibility scenarios.
+As defense technologies evolve and multi-target scenarios become increasingly prevalent, the demand for robust and accurate detection systems has intensified dramatically.
+I-A Challenges on Clustered Infrared Small Target Detection
+Traditional infrared small target detection research has predominantly focused on sparse target scenarios [1, 2], where isolated objects are distributed across relatively uniform backgrounds.
+However, contemporary real-world applications present fundamentally different challenges.
+In defense against drone swarms, multi-target tracking systems, and large-scale maritime rescue operations, targets frequently appear in clustered configurations, creating dense distributions that fundamentally alter the detection paradigm.
+This gap underscores an urgent need for advanced technology that can effectively identify and track clustered infrared small targets.
+This transition from sparse to dense target detection introduces a cascade of technical challenges that render existing methodologies inadequate:
+  1. 1.
+    Methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies.
+  2. 2.
+    Local contrast based methods [3], which depend on detecting objects against uniform backgrounds, become unreliable as dense clusters of targets blur the line between target and background.
+  3. 3.
+    Even state-of-the-art deep learning methods [4, 5, 6] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, adding to the complexity of detection tasks.
+These challenges are compounded by the limitations in current datasets, which primarily focus on sparse target detection and fail to address the complexities of clustered targets.
+I-B Limitations of Current Approaches
 While the rapid advancements in generic object detection have led to state-of-the-art methods like TOOD [7] and CrossKD [8] that address dense object detection, these techniques cannot be directly applied to the unique challenges posed by infrared small target detection.
-The specialized nature of this field presents several critical factors that hinder the direct transplantation of generic vision approaches:
+The specialized nature of this field presents two critical factors that hinder the direct transp lantation of generic vision approaches:
   1. 1.
-    Scarcity of Intrinsic Target Features: The foremost challenge stems from the paucity of intrinsic target features.
-Unlike generic objects, which boast rich visual cues such as appearance, texture, and structure, infrared small targets are minuscule in size, merely popping out as nondescript blobs in the image.
-This “visual poverty” renders recognition and localization methods based on target-specific characteristics largely ineffectual.
+    Scarcity of Target-Specific Features: Infrared small targets typically occupy only a few pixels and lack distinctive visual characteristics such as texture, shape, or structural details.
+This “visual poverty” renders traditional feature based recognition and localization methods largely ineffective, as there are insufficient discriminative cues to differentiate targets from background elements.
   2. 2.
-    Abundance of Similar Background Distractors: Infrared imagery is often plagued by a multitude of background distractors that bear a striking resemblance to the targets of interest.
-These distractors, such as noise spots and hot spots, also manifest as blobs, making it nearly impossible to discriminate between them and genuine targets based on local visual comparisons.
-Liao et al.
-[14] tackled similar challenges in RGB-thermal salient object detection, where distortions like noise and motion blur compounded the complexity of separating salient objects from background clutter.
+    Abundance of Similar Background Distractors: Infrared imagery frequently contains background elements [9], such as thermal noise, hot spots, and structural artifacts, that exhibit visual similarity to genuine targets.
+This similarity creates a high false alarm rate when relying solely on local appearance-based discrimination.
 These challenges inevitably lead us to ponder a fundamental question: What truly distinguishes visually similar objects in single-frame infrared small target detection, assigning them different class labels?
 Unraveling this enigma holds the key to developing effective and robust algorithms for this task.
-I-B Motivation
-In this paper, we posit that the answer to the aforementioned question lies in the background semantics.
-As illustrated in Fig.
-1, visually similar small objects at a distance can be classified differently based on their surrounding context.
-A small object in the sky background is likely to be a target of interest, while the same object against a building backdrop might be a false alarm.
-This stark contrast underscores the critical role that background semantics play in assigning meaningful labels to visually similar objects.
+I-C Motivation
+The answer to the aforementioned challenge lies in a paradigm shift: rather than relying solely on target appearance, we propose to exploit background semantics.
+Consider the scenario illustrated in Fig.
+1: when examining only local regions, distinguishing between genuine targets and false alarms becomes nearly impossible due to their visual similarity.
+However, by incorporating global contextual information, the classification task becomes significantly more tractable.
+A small bright object appearing against a sky background typically represents a target of interest, while an identical object against building infrastructure is more likely a false alarm.
+This contextual disambiguation demonstrates that background semantics[10] serve as crucial semantic anchors that enable accurate target identification where local appearance fails.
 However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research.
 Despite the availability of several public datasets [11, 12, 13, 5], none of these datasets provide detailed semantic annotations for the background regions.
-To enable the explicit semantic modeling of backgrounds, a prerequisite is the creation of a dataset that includes background class annotations.
-Such a dataset would provide a solid foundation for researchers to develop algorithms that can leverage the rich contextual information present in the background regions.
-To address the aforementioned challenges, this study presents DenseSIRST, a new open-source infrared target detection benchmark dataset.
-Compared to previous datasets in this field, DenseSIRST has implemented two significant upgrades.
-Firstly, the dataset transitions from a simple sparse target detection task to the precise detection of dense, clustered targets, substantially increasing the task’s difficulty and practical value.
-Secondly, DenseSIRST performs detailed per-pixel semantic segmentation of the background, dividing it into sky and non-sky categories, thereby providing preliminary support for explicit semantic modeling of the background.
-Building upon this foundation, we propose the Background-Aware Feature Exchange Network (BAFE-Net) framework, which transforms the infrared small target detection paradigm from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation.
-BAFE-Net introduces a dynamic cross-task feature hard-exchange module in the target detection head and background segmentation head, which selectively embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter.
-To the best of our knowledge, this paper is the first to propose a cross-task feature hard-exchange mechanism.
-Benefiting from the explicit modeling of background semantics, BAFE-Net can better understand the contextual information in the scene, thereby significantly improving the accuracy of target detection and greatly reducing the false alarm rate.
-Finally, by leveraging the fine-grained semantic annotations of the background, we propose a Background-Aware Gaussian Copy-Paste (BAG-CP) method.
-During the training phase, this method selectively pastes small targets into the sky region based on different background environments during data augmentation.
-This approach avoids pasting small targets into complex non-sky background regions, such as urban ground, which can produce highly similar false alarm targets.
+Without these annotations, it is challenging to explicitly model the background semantics necessary for improving detection performance.
+To bridge this critical gap and validate our hypothesis, we introduce DenseSIRST.
+This is a comprehensive infrared target detection benchmark that addresses both the clustered target detection challenge and the background semantic modeling requirement.
+Our dataset implements two revolutionary improvements over existing resources:
+  • Dense Target Scenarios: Transitioning from sparse to clustered target configurations that reflect real-world deployment conditions such as drone swarm detection and multi-target tracking applications.
+  • Background Semantic Annotations: Providing detailed per-pixel semantic segmentation of background regions, categorized into sky and non-sky classes, thereby enabling explicit background semantic modeling for the first time in this domain.
+Building upon the semantic-rich DenseSIRST dataset, we propose the Background-Aware Feature Exchange Network(BAFE-Net) framework, which fundamentally transforms infrared small target detection from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation.
+The core innovation of BAFE-Net lies in its dynamic cross-task feature hard-exchange module, which facilitates selective semantic information transfer between detection and segmentation tasks through an adaptive feature adapter.
+This represents the first implementation of cross-task feature hard-exchange mechanisms in infrared target detection.
+By explicitly modeling background semantics, BAFE-Net achieves superior contextual understanding, resulting in significant improvements in detection accuracy while substantially reducing false alarm rates.
 In summary, our contributions can be categorized into three main aspects:
   1. 1.
-    We propose a new task—-clustered infrared small target detection, and have released a new dataset, DenseSIRST, annotated with background semantic segmentation.
+    We propose a new task—-clustered infrared small target detection, which poses unique challenges compared to traditional single-target detection.
   2. 2.
-    Utilizing DenseSIRST, we propose and demonstrate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which outperforms the original Copy-Paste method that ignores background semantics.
+    We have released a new dataset, DenseSIRST, annotated with background semantic segmentation. This dataset offers rich contextual information to enhance target detection.
   3. 3.
     We propose the BAFE-Net, which achieves a dynamic mechanism of mutual promotion between segmentation and detection tasks by implementing selective feature hard-exchange with a cross-task adapter, based on explicit background modeling.
-II Related Work
-II-A Infrared Small Target Detection
-Infrared small target detection methods can be broadly classified into two main categories: model-driven and data-driven approaches.
-The model-driven methods, in particular, encompass three distinct subclasses, each employing unique strategies to capture and represent the relationship between targets and their surrounding backgrounds.
-  1. 1.
-    Background Estimation Methods: By leveraging the local consistency of backgrounds, these methods construct filters to enhance target saliency.
-However, they falter when dealing with extremely weak targets, leading to missed detections and a high sensitivity to hyperparameters.
-The limitation stems from insufficient semantic understanding of the features [17], which results in poor performance in complex backgrounds and increased false alarms.
-  2. 2.
-    Human Visual System (HVS) Methods: Techniques such as local contrast measure (LCM) [6] and multi-scale patch-based contrast measure (MPCM) [18] attempt to emulate the high spatial local contrast sensitivity of the human eye.
-Despite their intuitive appeal, these methods struggle with very small or blurred targets, mainly due to their inability to generate sufficient contrast.
-  3. 3.
-    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image (IPI) model [4] and reweighted infrared patch-tensor (RIPT) model [5].
-These methods have gained popularity due to their superior background suppression capabilities compared to previous techniques.
-However, they severely suffer from the challenge of distinguishing background distractors from targets due to the similar sparse characteristic.
-Despite their contributions, these methods suffer from several limitations, including poor performance in complex backgrounds, high false alarm rates, and sensitivity to hyperparameters, which we attribute to the lack of feature semantics.
-Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging deep networks to automatically learn features relevant for detecting small targets.
-Recent advancements have also focused on implementing these deep learning-based methods efficiently, such as the work on high-accuracy and real-time two-stage small object detection on FPGA [19].
-Methods such as asymmetric context modulation (ACM) [15], dense nested attention networks (DNANet) [16], and gated shaped trans U-Net (GSTUnet) [9] have shown significant promise.
-These methods concentrate on integrating the features of infrared small targets with neural networks, utilizing deeper neural networks or more robust attention modules.
-Similar advancements in applying deep learning to infrared imaging have been seen in related fields, such as the work on infrared-guided eye localization for autostereoscopic displays [20].
-The advantages of data-driven methods lie in their ability to automatically extract features and focus on multi-scale feature interaction, which can be seen as a form of implicit modeling of the relationship between targets and background.
-Similar approaches have been applied in other domains, such as cloud detection, where multi-level information fusion with edge information injection has been used to improve detection performance [21].
-In a related field, transformer-based models have been employed for spatio-temporal unsupervised traffic anomaly detection in aerial videos [22], demonstrating the versatility of these advanced deep learning techniques in handling complex visual data with small objects or anomalies.
-While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances.
-Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance [23].
-Their performance is greatly affected in dense scenes with multiple adjacent targets, where the assumption of local contrast between targets and their neighboring background no longer holds.
-This fundamental shift in target distribution necessitates the development of approaches that can effectively handle the complex interactions and dependencies among multiple targets in close proximity.
-In contrast to previous studies, our work explores a new detection paradigm for infrared small targets:
-  1. 1.
-    Unexplored Challenge for Clustered Targets: Unlike traditional approaches that focus on sparse targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field.
-We further facilitate the research by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
-  2. 2.
-    Explicit Semantic Modeling: In contrast to previous studies that treat backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic annotations for background modeling.
-Our proposed BAFE-Net thereby gains an enhanced understanding of the target’s background, leading to improved detection accuracy.
-  3. 3.
-    Introduction of Cross-Task Feature Exchange: Our dynamic cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not found in existing work.
-This approach enables a more precise differentiation between targets and their backgrounds.
-II-B Data Augmentation for Object Detection
-Data augmentation is pivotal in enhancing the performance and generalization of object detection models.
-By diversifying training data through various transformations such as flipping, rotation, scaling, and cropping, models learn to recognize and generalize across a broader spectrum of object variations.
-However, while these basic techniques are widely used, they often fall short in complex detection scenarios, necessitating more sophisticated methods.
-To address the limitations of traditional methods, advanced augmentation techniques have been developed, which introduce more nuanced transformations that better preserve and highlight essential features of the objects while still enhancing the dataset’s diversity.
-For instance, CutMix [24] improves upon the CutOut [25] method by overlaying a segment of one image onto another, blending labels proportionally to the area of the patch, thereby maintaining more contextual integrity.
-Attentive-CutMix [26] refines this approach by selecting regions based on feature importance, thus preserving significant target details.
-SnapMix [27] and SaliencyMix [28] adjust the blending process based on saliency or importance maps, aiming to preserve crucial features while enhancing background complexity.
-TransMix [29] integrates attention mechanisms to better manage the blending of features, focusing on maintaining important target characteristics during augmentation.
-StyleMix [30] and PuzzleMix [31] introduce style and structural considerations into the mix, allowing for more complex transformations that consider both content and style elements of the images, potentially offering more robust augmentation strategies.
-Despite these advancements, these methods primarily assume the presence of larger, easily distinguishable targets and do not specifically cater to the unique challenges posed by clustered infrared small target detection.
-None of these methods are explicitly designed to handle the high-density and occlusion-prone nature of infrared small targets.
-This gap highlights the need for augmentation strategies that directly address the intricacies of detecting small, closely-spaced targets against complex backgrounds.
-Although our BAG-CP aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and strategies:
-  1. 1.
-    Pioneering Data Augmentation for Clustered Infrared Small Target Detection: Uniquely, our method pioneers a data augmentation technique specifically tailored for clustered infrared small target detection.
-It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not considered in existing techniques.
-  2. 2.
-    Background-Aware Target Placement: Our BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics.
-Instead, it integrates background semantics to guide target placement, thereby skirting potential category conflicts induced by inconsistent background contextual semantics.
-II-C Multi-Task Learning and Cross-Task Mechanism
-Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning shared representations from related tasks, such as detection and segmentation.
-Takikawa et al.
-[32] introduced a two-stream CNN architecture for semantic segmentation that explicitly wires shape information as a separate processing branch, using gates to connect the intermediate layers of the two streams.
-Misra et al.
-[33] proposed a new sharing unit called “cross-stitch” unit, which combines activations from multiple networks and can be trained end-to-end to learn an optimal combination of shared and task-specific representations.
-Zhang et al.
-[10] developed ISNet, which employs a Taylor finite difference inspired edge block and two-orientation attention aggregation block to enhance edge information and capture shape characteristics of infrared small targets.
-However, these methods may not be well-suited for clustered infrared small target detection, as they often rely on edge detection or surface normal estimation, which assume larger targets.
-In real-world scenarios, infrared small targets typically consist of only 1 to 9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
-In this context, cross-task mechanisms that enable semantic exchange between different tasks are more relevant to our work.
-These methods enhance MTL by facilitating feature exchange or distillation across different tasks, allowing each task to benefit from complementary signals of others.
-Methods like prediction-and-distillation network utilize attention-based mechanisms to distill multi-modal features [34], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [35].
-While our research shares the underpinnings of multi-task learning and cross-task mechanisms, we pioneer a new approach to feature fusion.
-Instead of following the well-trodden path of soft fusion found in attention-based methods, our work introduces an innovative protocol–-a dynamic hard-exchange mechanism.
-To the best of our knowledge, this is the first work to explore the potential of dynamic hard-exchange of features between tasks.
-III The DenseSIRST Dataset
-Figure 2: Illustration of selected images from our proposed DenseSIRST dataset.
-Each image pair consists of a simulated dense small target image (top) and its corresponding sky-segmented version (bottom).
-These images highlight the challenges posed by densely clustered targets within diverse backgrounds.
-III-A Motivation and Rationale
-The detection of small, low-contrast targets against complex backgrounds in infrared imagery poses a significant challenge.
-Existing datasets, though instrumental in progressing the field, exhibit critical deficiencies.
-As shown in Table I, well-regarded datasets such as SIRST V1 [15], SIRST V2 [3], IRSTD1K [10], and SIRSTAUG [36] predominantly feature sparsely distributed targets and lack per-pixel background annotations.
-This stands in stark contrast to the dense, diminutive targets frequently encountered in real-world scenarios.
-To bridge this gap, we present DenseSIRST: a meticulously annotated dataset of infrared images teeming with densely clustered, small targets, which is a more realistic and challenging benchmark for algorithm training and evaluation.
-Fig.
-2 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations.
-These annotations provide valuable contextual information for developing and evaluating detection algorithms.
-TABLE I: Comparison of the DenseSIRST dataset with other popular datasets used in infrared small target detection.
+TABLE I: A Comparative Analysis of the DenseSIRST Dataset and Other Popular Datasets for Infrared Small Target Detection.
 Datasets Image Type Annotation Type Image Number Target Number Average Target Area Sparse or Clustered Background Semantic Annotation
@@
 DenseSIRST (Ours) Real + Synthetic Pixel + BBox + Point 1024 13655 6 Cluster
+II Related Work
+The core idea of traditional infrared small target detection algorithms is to suppress background interference and enhance target saliency. These methods mainly fall into three categories:
+  1. 1.
+    Background Estimation Methods: By exploiting the local smoothness or non-local self-similarity of background regions, various filtering strategies are designed to highlight targets.
+However, these methods often fail when facing extremely weak or low-contrast targets, primarily due to their limited semantic perception capabilities and high sensitivity to handcrafted hyperparameters.
+Consequently, they are prone to both missed detections and false alarms in complex backgrounds.
+  2. 2.
+    Human Visual System (HVS) Methods: Techniques such as local contrast measure (LCM) [3] and multi-scale patch-based contrast measure (MPCM) [15] attempt to emulate the high spatial local contrast sensitivity of the human eye.
+While intuitive and computationally efficient, their reliance on pixel-level or patch-level contrast leads to diminished performance when the target is extremely small, blurred, or embedded in cluttered textures.
+  3. 3.
+    Low-Rank and Sparse Decomposition Methods: Inspired by matrix and tensor decomposition theories, methods such as the Infrared Patch-Image (IPI) model [1], the Reweighted Infrared Patch-Tensor (RIPT) model [2] and the low-rank tensor approximation with saliency prior (LRTA-SP) model [16] decompose the image into low-rank background and sparse target components.
+Despite demonstrating strong background suppression capabilities, they struggle to differentiate true targets from structured background noise or distractors due to the shared sparsity characteristics, limiting their robustness in real-world scenarios.
+With the advent of deep learning, data-driven methods have significantly advanced infrared small target detection by learning robust representations directly from data, rather than relying on handcrafted priors.
+These approaches leverage deep neural networks to automatically extract semantic features, enabling better adaptation to diverse and complex backgrounds.
+Several representative methods have demonstrated the advantages of deep architectures in modeling contextual information and capturing multi-scale cues.
+For example, the asymmetric context modulation network (ACM) [11] introduces a context-sensitive structure to enhance the discrimination between targets and background.
+DNANet [13] further incorporates dense attention modules to refine features hierarchically, boosting detection performance across varied scenes.
+Recently, FATCNet [17] employs a dual-branch CNN-Transformer with adaptive fusion for effective local-global feature integration.
+Similarly, ILNet [18] enhances detection by emphasizing low-level salient features via lightweight fusion and dynamic aggregation.
+Moreover, MNHU [19] introduces a high-order UNet with nested multi-scale fusion to capture long-range dependencies and hierarchical representations.
+In addition, SFCANet [20] combines spatial-frequency attention with deep supervised multi-task learning to improve background suppression and target discrimination.
+While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances.
+This simplification limits the ability of the model to exploit contextual cues embedded in semantically distinct regions such as sky, sea, or land, which often exhibit different noise characteristics and target distributions.
+Moreover, although deep learning-based methods no longer explicitly rely on sparse priors or local contrast assumptions as traditional approaches do, many still implicitly incorporate such prior knowledge to improve detection performance.
+However, in dense scenes characterized by multiple closely spaced targets, the assumption of local contrast between a target and its surrounding background becomes invalid.
+This shift in spatial distribution not only weakens the effectiveness of contrast-based cues but also introduces significant ambiguity in distinguishing adjacent targets.
+Consequently, there is a pressing need for methods capable of modeling the complex interactions and contextual dependencies among targets in densely populated scenarios.
+To address these limitations, our work explores a novel detection paradigm for infrared small targets, focusing on two main aspects:
+  1. 1.
+    Unexplored Challenge for Clustered Targets: Departing from conventional assumptions of sparsity, we target the more realistic and challenging scenario of detecting densely clustered targets.
+To support this direction, we construct DenseSIRST, a specialized dataset designed to benchmark performance under such conditions.
+  2. 2.
+    Explicit Semantic Modeling: Rather than treating all backgrounds uniformly, we utilize the semantic annotations provided by DenseSIRST to enable fine-grained background-aware detection.
+Our proposed BAFE-Net leverages this information to enhance contextual understanding, thereby improving the discrimination between targets and background distractors.
+III The DenseSIRST Dataset
+Figure 2: Synthetic generation pipeline and representative samples of the DenseSIRST dataset.
+(a) The BAG-CP pipeline for synthesizing the DenseSIRST dataset incorporates three key components: semantic-aware background selection, Gaussian-based target fusion, and realistic sample generation.
+(b) Illustration of selected images from our proposed DenseSIRST dataset.
+Each image pair consists of a simulated dense small target image (left) and its corresponding sky-segmented version (right).
+Infrared small target detection is crucial in surveillance, early warning, and remote sensing.
+However, detecting small, low-contrast targets against complex and cluttered backgrounds remains a significant challenge.
+Existing benchmark datasets, such as SIRST V1 [11], SIRST V2 [12], IRSTD1K [5], and SIRSTAUG [14], have contributed substantially to algorithmic advancement, yet they predominantly feature sparsely distributed targets and lack comprehensive pixel-level background annotations, as demonstrated in Table I.
+The mismatch between current benchmarks and real-world scenarios hinders the development of methods for detecting dense, low-contrast targets within complex and semantically rich backgrounds.
+To address this gap, we present DenseSIRST, a novel infrared dataset tailored for dense infrared small target detection under complex backgrounds.
+DenseSIRST offers high-density target distributions and fine-grained pixel-level annotations, as illustrated in Fig.
+2 (b).
+By providing richer spatial and semantic context, it aims to foster the development of more robust and practically applicable detection algorithms.
+III-A Dataset Construction
+The DenseSIRST dataset is constructed by integrating synthetically generated small targets into infrared background images from the publicly available SIRST V2 [12] dataset.
+To achieve realistic and contextually consistent fusion, we propose a novel strategy called Background-Aware Gaussian Copy-Paste (BAG-CP), which builds upon the principles of mixup [21].
+Unlike naive copy paste, BAG-CP selectively pastes targets into semantically regions informed by background context, thereby enhancing the naturalness and variability of synthesized samples.
+The overall framework is depicted in Fig.
+2 (a).
+III-A1 Infrared Small Target Library Construction
+To simulate diverse and realistic small targets, we first build a target library by extracting individual small target patches from infrared datasets.
+Each extracted target is carefully cropped to include the target region with minimal background.
+This target library provides a rich source of varied target appearances, shapes, and sizes for subsequent synthesis.
+III-A2 Semantic-Aware Target Paste Region Selection
+To ensure plausible and contextually consistent target placement, the BAG-CP strategy incorporates semantic-aware selection of paste regions.
+Specifically, from each background image from SIRST V2, one to three candidate regions are randomly defined within relatively smooth and homogeneous sky areas.
+Each selected region is fixed to a size of $20\times 20$ pixels and serves as a dense paste area for simulating clustered target distributions.
+III-A3 Target Pasting and Fusion Process
+Within each selected region, 6 to 12 targets are randomly sampled from the target library and pasted with controlled scale and spacing.
+Each target is resized to a dimension no larger than $5\times 5$ pixels to mimic realistic scale variation.
+Targets are spaced with 1 to 2 pixels apart to emulate natural cluster formations typical in dense small target scenarios.
+To seamlessly blend the pasted targets with the background, an adaptive Gaussian fusion matrix $G$ is applied, producing smooth transitions at target edges and avoiding harsh visual boundaries.
+The synthesized target patch $T_{new}$ is computed as:
+$T_{new}=B_{raw}+T_{raw}\times\lambda\times G\ ,\lambda\in[0.5,1],$ (1)
+where $B_{raw}$ is the original background patch, $T_{raw}$ is the raw target instance, and $\lambda$ is a random brightness scaling factor introducing intensity diversity.
+The Gaussian matrix $G$ is defined by
+$G(x,y)=e^{-\frac{{(x-\rho_{x}w)^{2}}}{{2\sigma_{x}^{2}}}-\frac{{(y-\rho_{y}h)^{2}}}{{2\sigma_{y}^{2}}}},$ (2)
+where $w$ and $h$ represent the width and height of the target, respectively.
+The parameters $\rho_{x}$ and $\rho_{y}$ specify the Gaussian center offset relative to the target dimensions, randomly sampled from $[0,0.2]$.
+The standard deviations $\sigma_{x}$ and $\sigma_{y}$, controlling the spread of the Gaussian kernel, are independently sampled from $[0.3,0.6]$, which enables varied edge softness and thus improves the naturalness of target boundaries.
+III-A4 Geometric and Appearance Augmentation
+To further increase target diversity, each pasted target undergoes a random rotation transformation with angle $\theta$ uniformly sampled from $90^{\circ},90^{\circ}]$, implemented by
+$\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x-\mu_{x}\\ y-\mu_{y}\end{bmatrix}$, (3)
+where $(\mu_{x},\mu_{y})$ is the center of the target patch.
 Figure 3: Statistical characteristics of our DenseSIRST dataset, underscoring the detection challenges it presents.
@@
 (c) The brightness distribution of small targets in the dataset, highlighting the wide spectrum from dark to light targets this dataset offers for detection tasks.
-III-B Target Synthesis and Composition
-The DenseSIRST dataset, a blend of real and synthetic infrared images, provides a challenging and dense environment for training robust detection algorithms.
-We curated this dataset from diverse real-world scenarios, supplemented with synthetic targets designed for physical plausibility and visual realism.
-The dataset leverages our SIRST V2 dataset, creating a broad library of infrared small targets.
-Each image is meticulously designed to emulate the densely populated surveillance scenarios.
-Specifically, one to three “dense areas” are defined randomly within the less cluttered sky regions of each base image.
-These areas, each measuring $20\times 20$ pixels, are populated with 8 to 12 targets selected randomly from the target library.
-The selected targets are resized to dimensions within $5\times 5$ pixels to maintain scale variability and pasted with a spacing of 1 to 2 pixels between them to simulate realistic clustering.
-To ensure a smooth integration of targets and their backgrounds, we employ a Gaussian-weighted copy-paste technique, an evolution of the concept from mixup.
-This modification is governed by the following equation:
-$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1],$ (1)
-where $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location.
-Different from mixup, its weight is set to $1$ for enhancing the reality of the synthesized target.
-$B_{box}$ represents the original target appearance.
-The scalar $\lambda$, which adjusts the target’s brightness, is randomized to diversify the brightness and contrast of targets, thus enriching the sample space.
-The matrix $G$ denotes a Gaussian blur, employed during target pasting, and is defined as:
-$G(x,y)=e^{-\frac{{(x-\rho_{x}w)^{2}}}{{2\sigma_{x}^{2}}}-\frac{{(y-\rho_{y}h)^{2}}}{{2\sigma_{y}^{2}}}},$ (2)
-where the relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$.
-These values, when multiplied by the target bounding box’s width $w$ and height $h$, yield the actual pixel offsets, ensuring appropriate scaling with the target size.
-The Gaussian distribution’s spread is controlled by the standard deviations $\sigma_{x}$ and $\sigma_{y}$ along the $x$ and $y$ axes, randomized within $[0.3,0.6]$.
-This randomness guarantees varying degrees of edge blurring among the simulated targets, reinforcing their realistic appearance.
-To further enhance variability, each target undergoes a rotational transformation.
-The Gaussian matrix’s rotation $\theta$ is applied as:
-$\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x-\mu_{x}\\ y-\mu_{y}\end{bmatrix},$ (3)
-where $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
-III-C Dataset Description and Characteristics
-III-C1 Dataset Composition
+III-B Dataset Description and Characteristics
+III-B1 Dataset Composition
 The DenseSIRST dataset comprises 1024 infrared images, with a total of 13,655 densely clustered small targets.
@@
 The division of training, validation, and test set is aligned with the SIRST V2 dataset.
-III-C2 Annotations and Labels
+III-B2 Annotations and Labels
 In the DenseSIRST dataset, each image is meticulously annotated with three types of labels: pixel-level masks, bounding boxes, and point annotations.
@@
 Point annotations mark the center of each target, allowing for the evaluation of localization accuracy.
-III-C3 Statistical Characteristics
-The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection. This part offers an analysis of the dataset’s key characteristics.
-Target Size Distribution: Most targets, as indicated in Fig.
-3 (b), are extremely small, predominantly below $5\times 5$ pixels.
-The $3\times 3$ pixels size is the most frequent, aligning with the characteristics of small target detection in infrared imagery.
-Local Contrast Distribution: The 90% targets have local contrast below $2$, as shown in Fig.
-3 (c).
+III-B3 Density Definition
+To quantitatively capture the degree of target clustering, we define a density measure tailored for cluster regions within the dataset.
+A cluster region is identified when the Euclidean distance between the centroids of any two targets is less than 10 pixels, representing a meaningful spatial proximity threshold for dense grouping.
+Formally, the density $\mathcal{D}$ of a cluster region is computed as:
+$\mathcal{D}=\mathcal{N}\times\frac{\mathcal{A}_{t}}{\mathcal{A}_{c}},$ (4)
+where $\mathcal{N}$ denotes the number of targets within the cluster, $\mathcal{A}{t}$ is the cumulative pixel area covered by all target masks inside the cluster, and $\mathcal{A}{c}$ represents the total pixel area of the cluster region itself.
+This metric jointly reflects both the quantity of targets and their relative spatial occupancy, providing a nuanced descriptor of local target density.
+For consistent comparison across images, density values are normalized via min-max scaling based on the global minimum and maximum densities observed in both clustered and isolated target regions.
+Empirical statistics reveal a pronounced density difference, cluster regions exhibit an average density of 0.4092 compared to 0.1342 in single target regions, underscoring the metrics effectiveness in distinguishing densely grouped targets from sparse ones.
+III-B4 Statistical Characteristics
+Fig. 3 illustrates key statistical properties of the DenseSIRST dataset, emphasizing the detection challenges it entails.
+  • Target Size Distribution (Fig.
+3 (a)): The vast majority of targets are extremely small, predominantly under $5\times 5$ pixels, with the most frequent size being $3\times 3$ pixels.
+This distribution aligns with the typical scale of small infrared targets.
+  • Local Contrast Distribution (Fig.
+3 (b)): Approximately 90% of targets possess a local contrast below 2 relative to their surrounding background.
 This low contrast is indicative of the inherent difficulty in detecting such targets against cluttered backgrounds.
 Moreover, the presence of nearby targets within dense clusters can inadvertently be included in the background calculation, leading to an underestimation of local contrast and further complicating detection.
-Target Brightness Distribution: Fig.
-3 (d) reveals that only a minority of targets are the brightest points in their respective images.
-This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection.
-In conclusion, the DenseSIRST dataset, with its wide range of image dimensions, prevalence of small target sizes, and low local contrast, provides a realistic and challenging benchmark for infrared small target detection algorithms.
+  • Target Brightness Distribution (Fig.
+3 (c)): The brightness values of targets span a wide spectrum, with only a minority being the brightest points in their respective images.
+This diversity invalidates brightness as a reliable standalone feature, highlighting the necessity for robust algorithms that leverage richer contextual and structural information.
+In conclusion, the DenseSIRST dataset exhibits rich annotation modalities, diverse environmental scenarios, and statistically challenging target characteristics.
+These attributes collectively establish it as a comprehensive and realistic benchmark, particularly well suited for developing and evaluating infrared small target detection algorithms under dense and low-contrast conditions.
 IV Background-Aware Feature Exchange Network
 IV-A Overall Architecture and Training Pipeline
-Figure 4: Schematic diagram of the proposed BAFE-Net network. BAFE-Net contains three modules: ResNet, FPN and BAFE-Head.
-Figure 5: First, a dense area is randomly selected from the sky area of the image, and then several small targets are randomly selected from the small target library.
-After determining the appropriate pasting position, the Gaussian technique is used to integrate the small targets into the new background.
-The overall architecture of BAFE-Net, which adopts the fully convolutional one-stage object detector (FCOS) [37] as its host network, is illustrated in Fig.
+Figure 4: The framework of BAFE-Net.
+BAFE-Net contains three modules: ResNet, FPN and BAFE-Head.
+The Head module integrates a background segmentation branch to work in parallel with the classification branch.
+Both branches utilize the DCS Module to select the most discriminative top-k channel features, followed by channel-wise feature interaction.
+This enhancement transforms the original single-task object detection framework into a multi-task learning architecture, enabling simultaneous and efficient execution of both object detection and background semantic segmentation.
+The proposed BAFE-Net architecture, which builds upon the fully convolutional one-stage object detector (FCOS) [22], is illustrated in Fig.
 4.
-In addition to the original detection head present in the baseline, BAFE-Net introduces a segmentation head within the BAFE-Head module.
-The primary purpose of incorporating this segmentation head is to enable explicit modeling of background semantics, which is crucial for enhancing the network’s ability to distinguish between real targets and background clutter in infrared scenes.
-By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information.
-Furthermore, BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to facilitate the bidirectional flow of information between the detection and segmentation branches.
+The network comprises three key components: ResNet backbone, Feature Pyramid Network (FPN), and BAFE-Head module.
+The BAFE-Head features two parallel branches: the original FCOS detection branch and an innovative background segmentation branch.
+This dual-branch design enables explicit modeling of background semantics, which is essential for distinguishing real targets from background clutter in infrared scenes.
+BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism that facilitates bidirectional information flow between the detection and segmentation branches.
+Both branches incorporate a Dynamic Channel Selection (DCS) module, which identifies and utilizes the top-k most discriminative channel features to enable efficient Channel-wise Feature Exchange.
 This dynamic exchange allows the network to adaptively leverage the most relevant features from each task, leading to improved detection performance.
-To address the challenge of limited target samples, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP).
-This technique strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, ensuring that the pasted targets are only placed in sky regions.
-The use of Gaussian smoothing at the edges of the pasted targets ensures seamless integration, avoiding artificial boundary artifacts that could mislead the learning algorithm.
-By augmenting the training dataset with realistically integrated synthetic targets, BAG-CP allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds.
+To address the challenge of limited target samples, BAFE-Net adopts the dataset synthesis method BAG-CP as data augmentation strategy.
+By leveraging semantic context, this method pastes synthetic infrared small targets exclusively onto sky regions in real-scene backgrounds, with Gaussian smoothing applied to the edges for seamless integration.
 This context-aware synthesis and integration approach enables BAFE-Net to effectively recognize and handle these challenging conditions, ultimately setting a new standard for performance in infrared small target detection.
 IV-B Dynamic Cross-Task Feature Exchange Head
-For infrared small target detection, the scarcity of intrinsic target features and the abundance of similar background distractors pose significant challenges.
-To address this, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation.
-The motivation behind this cross-task feature exchange is twofold:
-  1. 1.
-    Complementary Nature of Tasks: Although target detection and background segmentation have different objectives, they share a common goal of understanding the scene.
-By exchanging features between these tasks, each task can benefit from the complementary information learned by the other, leading to a more comprehensive understanding of the image.
-  2. 2.
-    Synergistic Relationship: Target detection focuses on identifying the presence and location of small targets, while background segmentation emphasizes the contextual information and local details of the scene.
-By sharing features between these tasks, target detection can leverage the semantic understanding of the background to better distinguish targets from distractors, while background segmentation can utilize the localized target information to refine its segmentation boundaries.
-IV-B1 Dynamic Channel Selection
-To further optimize cross-task feature exchange, we propose a dynamic selection mechanism that determines the most relevant feature channels for exchange based on their importance to task performance.
+Infrared small target detection presents inherent challenges due to the weak discriminative characteristics of targets and the prevalence of complex background clutter.
+Conventional single-branch detection frameworks often struggle to effectively distinguish true targets from background distractors, particularly in scenarios characterized by limited target information and ambiguous contextual cues.
+To address these challenges, we propose a multi-task collaborative framework that integrates object detection and background semantic segmentation within a unified prediction head, referred to as BAFE-Head.
+The core motivation behind this design lies in the complementary nature of the two tasks: object detection is tailored to localize small, sparse targets, whereas semantic segmentation captures global scene semantics and background structures.
+By jointly optimizing these tasks, the proposed framework leverages background semantics to suppress false positives and utilizes target localization cues to refine segmentation boundaries.
+This joint formulation facilitates a more holistic and robust feature representation, enhancing the discriminative capability of the network in cluttered infrared scenes.
+IV-B1 Dynamic Channel Selection Module
+Unlike conventional multi-task learning strategies that typically employ static or full feature sharing, which may lead to the propagation of redundant or even conflicting information across tasks, we introduce a dynamic cross-task feature exchange mechanism.
+This mechanism adaptively selects and exchanges task-relevant channel-wise features between the detection and segmentation branches.
 The importance of each channel is estimated using an attention mechanism, which assigns weights to channels according to their contribution.
-Given an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature map, respectively, we first compute a channel-wise attention weight vector $\mathbf{m}\in\mathbb{R}^{C}$ by a mechanism inspired by the design of our previous work Large Selective Kernel Network (LSKNet) [38].
+This dynamic and selective interaction not only strengthens the synergy between detection and segmentation but also effectively mitigates the risk of negative transfer, thereby enhancing overall detection performance in complex infrared environments.
+Projection Module.
+Given an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature map, respectively, we design a Projection Module to generate a channel-wise attention vector $\mathbf{m}\in\mathbb{R}^{C}$.
+This design is inspired by our previous work, Large Selective Kernel Network (LSKNet) [23], and aims to enhance channel discrimination by integrating local feature extraction with global context modeling.
 This is achieved by applying a series of operations to $\mathbf{X}$, including local feature extraction, channel-wise statistics computation, and weighted fusion.
-The process can be summarized as follows:
-$\mathbf{m}$ $=\sigma\left(\mathbf{X}\odot\left(\mathcal{F}(\mathbf{X})\odot\sigma\left(\mathcal{S}\left(\text{avg}(\mathbf{X})\oplus\max(\mathbf{X})\right)\right)\right)\right),$ (4)
-where $\odot$ denotes element-wise multiplication, $\oplus$ denotes the concatenation operation, and $\sigma$ is the Sigmoid function.
-$\max$ and avg are the maximum and average pooled spatial feature descriptors.
-$\mathcal{S}$ is a convolution operation that reduces the dimensionality of concatenated features.
-$\mathcal{F}$ is also a convolution operation applied to $\mathbf{X}$ , extracting local features.$\mathbf{m}\in\mathbb{R}^{C}$ is a one-dimensional tensor, and $C$ is the number of channels.
+  • local feature extraction: To capture fine-grained local representations, we first apply two independent convolutional operations to the input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, each reducing the channel dimension to $C/2$.
+These two branches are designed to extract complementary local patterns.
+    $\mathbf{X}_{1}$ $=\text{Conv}_{1}(\mathbf{X}),$ (5)
+    $\mathbf{X}_{2}$ $=\text{Conv}_{2}(\mathbf{X}),$
+    $\mathbf{X}_{c}$ $=\mathbf{X}_{1}\oplus\mathbf{X}_{2},$
+    where $\oplus$ denotes concatenation operation, resulting in $\mathbf{X}_{c}\in\mathbb{R}^{C\times H\times W}$.
+  • channel-wise statistics computation: To integrate global context and emphasize informative regions, we apply both average pooling and max pooling along the channel dimension to form a compact spatial descriptor, followed by a convolutional projection and sigmoid activation.
+    $\text{Feat}=\sigma(\mathcal{S}(\text{avg}(\mathbf{X}_{c})\oplus\max(\mathbf{X}_{c}))),$ (6)
+    where $\max$ and avg are the maximum and average pooled spatial feature descriptors.
+$\mathcal{S}$ is a convolution operation that reduces the dimensionality of concatenated features, and $\sigma(\cdot)$ is the sigmoid function.
+The obtained $\text{Feat}\in\mathbb{R}^{2\times H\times W}$ is a tensor, which represents the weighted fusion of $\max(\mathbf{X})$ and $\text{avg}(\mathbf{X})$.
+  • weighted fusion: The attention map Feat is split along the channel axis and used to reweight the previously extracted features $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ through element-wise multiplication.
+The reweighted features are then concatenated to form the output of the projection module:
+    $\mathbf{Z}$ $=\mathcal{F}(\mathbf{X}_{1},\text{Feat})\oplus\mathcal{F}(\mathbf{X}_{2},\text{Feat}),$ (7)
+    $\mathbf{m}$ $=\text{mean}\left(\sigma\left(\mathbf{X}\odot\mathbf{Z}\right)\right),$
+    where $\odot$ denotes element-wise multiplication, and the operator $\mathcal{F}(\mathbf{X}_{i},\text{Feat})$ denotes a channel-wise spatial reweighting operation that modulates the local feature $\mathbf{X}_{i}$ using the corresponding attention map $\text{Feat}_{i}$.
+mean denotes a global average pooling operation that aggregates feature representations across spatial dimensions, reducing spatial information into a channel-wise global descriptor.
+$\mathbf{m}\in\mathbb{R}^{C}$ is a one-dimensional tensor, and $C$ is the number of channels.
 Each element $\mathbf{m}$ represents the attention logits for the $i$-th channel.
@@
 where ${V}_{\text{topk}}\in\mathbb{R}^{K}$ represents the top $K$ largest weight values for each sample.
-${I}_{\text{topk}}\in\mathbb{R}^{K}$ represents the corresponding channel indices for these weights.
-This process dynamically selects the channels for exchange, improving the effectiveness of cross-task feature sharing.
-IV-B2 Cross-Task Feature Hard-Exchange
-Within BAFE-Net, we introduce a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks.
-Feature hard-exchange allows for the direct transfer of learned features from one task to another.
-By swapping entire feature channels, the receiving task has access to the explicit representation learned by the other task, facilitating a more direct integration of complementary information.
-To perform feature exchange, let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ be the input tensors from the target detection and background segmentation tasks, respectively, each with a shape of $(C,H,W)$, where $C$ is the number of channels, and $H$ and $W$ are the height and width.
-We define an exchange mask $\mathbf{e}\in\mathbb{R}^{C}$:
-$\mathbf{e}[i]=\begin{cases}1,&\mathrm{if}\ i\in I_{\text{topk}}\\ 0,&\mathrm{otherwise}\end{cases},$ (6)
-where $i$ represents the channel index. The exchange mask $\mathbf{e}$ is then expanded to the shape of $(C,H,W)$:
-$\mathbf{E}_{d}=\mathbf{e}\otimes\mathbf{1}_{H\times W},$ (7)
-where $\otimes$ denotes the expansion operation, and $\mathbf{1}_{H\times W}$ is a tensor of ones with a shape of $(H,W)$.
-$\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ can obtain $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ respectively through Eq.
-(4) to Eq.
-(7).
-The output tensors $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ after the hard feature exchange are computed as:
-$\mathbf{Y}_{1}=\mathbf{X}_{1}\odot(\mathbf{1}-\mathbf{E}_{d1})+\mathcal{A}\left(\mathbf{X}_{2}\odot\mathbf{E}_{d2}\right),$ (8)
-$\mathbf{Y}_{2}=\mathbf{X}_{2}\odot(\mathbf{1}-\mathbf{E}_{d2})+\mathcal{A}\left(\mathbf{X}_{1}\odot\mathbf{E}_{d1}\right),$ (9)
-where $\odot$ denotes element-wise multiplication, $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ represent the output features, and $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ are the dynamic masks.
-To address the cross-modal nature of features, we introduce an adapter module $\mathcal{A}$, implemented as a multi-layer perceptron (MLP), which projects features into a common space before exchange.
-This design enables effective cross-modal interaction while preserving modality-specific characteristics.
-IV-C Background-Aware Gaussian Copy-Paste
-Infrared small target detection is inherently challenged by the paucity of intrinsic visual features in targets and the interference from similar background elements.
-Traditional data augmentation strategies, such as random cropping or flipping, are not effective in addressing these unique challenges.
-Consequently, we propose a novel data augmentation technique, Background-Aware Gaussian Copy-Paste (BAG-CP), which capitalizes on contextual information to enhance the training process.
-BAG-CP selectively augments training images by pasting infrared small targets into semantically relevant background areas, generating more realistic and diverse training samples.
-The fundamental idea is to bolster the detector’s ability to discern real targets from background noise through the seamless blending of pasted targets into their new environment.
-As illustrated in Fig.
-5, The BAG-CP process unfolds as follows:
-Prerequisite: Semantic Annotation of Background: Before performing the copy-paste operation, it is crucial to have the image segmented into different semantic regions.
-The DenseSIRST dataset includes annotated background categories, providing detailed background masks that separate sky from non-sky regions.
-Step 1: Selection and Extraction of Infrared Small Targets: A library of small targets is created during the generation of the DenseSIRST dataset.
-A random selection of small targets from this library is made for pasting into the current training image.
-Step 2: Gaussian Smoothing of Edges: Prior to pasting, the selected targets are subjected to Gaussian smoothing.
-This step ensures a smooth transition around the pasted targets, minimizing any conspicuous artifacts that might disrupt the learning process.
-Step 3: Context-Aware Target Pasting Based on Background Semantics: The core of BAG-CP lies in context-aware placement.
-Using the semantic annotation of the background, we classify it into sky and non-sky regions.
-When choosing where to paste the Gaussian-smoothed targets, we only select locations within the sky regions.
-This approach mimics real-world scenarios where infrared small targets, especially aerial targets, are more likely to appear against a sky background.
-During the pasting process, 1-3 regions are selected within the image to form clusters of small targets.
-Each region is 20 × 20 pixels in size and is populated with 8 to 12 targets randomly chosen from the target library.
-The selected targets are resized to within 5 × 5 pixels to maintain proportional variability and are pasted with a spacing of 1 to 2 pixels to simulate realistic clustering.
-The Gaussian-weighted edge blurring technique used follows Eq.
-(1), Eq.
-(2), and Eq.
-(3).
-V Experiments
-In this section, we outline a series of experimental evaluations designed to validate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method and the Background-Aware Feature Exchange Network (BAFE-Net) proposed in this study.
-Additionally, we conduct comparative analyses with state-of-the-art methods in infrared small target detection, using both quantitative metrics and qualitative assessments.
-We address the following pivotal questions in our experimental analysis:
-  1. 1.
-    Ablation Study on BAG-CP. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
-  2. 2.
-    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlight the contributions of each component in BAFE-Net.
-  3. 3.
-    Feature Exchange Mechanism Analysis.
-We investigate the impact of different feature exchange strategies on the performance of BAFE-Net to determine the optimal approach for exchanging features between the target detection head and the background segmentation head.
-  4. 4.
-    Impact of Backbone Depth: We examine the influence of varying backbone depths on the performance of the BAFE-Net.
-The primary aim is to establish the optimal depth that maximizes target detection while balancing computational efficiency.
-  5. 5.
-    Comparative Analysis with State-of-the-Art Methods. We conduct a comprehensive comparison between our method and other state-of-the-art methods on the DenseSIRST dataset.
-V-A Experimental Settings
-1) Dataset: During the experimental phase of our research, we conducted tests on the proposed dataset DenseSIRST.
-2) Implementation Details: Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion.
-We choose Focal Loss as the loss function of the classification branch, IoU Loss as the loss function of the regression branch, and Cross Entropy Loss as the loss function of the segmentation branch.
-We choose DAdaptAdam as the optimizer, learning rate is set to 1.0, and weight decay is set to 0.05.
-Considering training and convergence time, we trained for 20 epochs within the DeepInfrared framework.
-3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment.
-TABLE II: Comparison with Other State-of-the-art methods on DenseSIRST.
-Method Backbone DenseSIRST
-FLOPs $\downarrow$ Params $\downarrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-              One-stage
-SSD [39] 87.552G 23.746M 0.211 0.421 0.178 0.424
-RetinaNet [40] ResNet50 52.203G 36.330M 0.114 0.510 0.086 0.523
-YOLOv3 [41] 50.002G 61.949M 0.233 0.424 0.207 0.413
-CenterNet [42] ResNet50 50.278G 32.111M 0.138 0.316 0.124 0.317
-FCOS [37] ResNet50 50.291G 32.113M 0.232 0.315 0.204 0.324
-ATSS [43] ResNet50 51.504G 32.113M 0.248 0.327 0.202 0.326
-CentripetalNet [44] 0.491T 0.206G 0.244 0.259 0.201 0.244
-AutoAssign [45] ResNet50 50.555G 36.244M 0.255 0.354 0.180 0.314
-GFL [46] ResNet50 52.296G 32.258M 0.264 0.367 0.230 0.317
-PAA [47] ResNet50 51.504G 32.113M 0.255 0.545 0.228 0.551
-VFNet [48] ResNet50 48.317G 32.709M 0.253 0.336 0.214 0.336
-PVT-T [49] 41.623G 21.325M 0.109 0.481 0.093 0.501
-YOLOF [50] ResNet50 25.076G 42.339M 0.091 0.009 0.002 0.009
-YOLOX [51] 8.578G 8.968M 0.210 0.341 0.180 0.331
-TOOD [12] ResNet50 50.456G 32.018M 0.256 0.355 0.226 0.342
-DyHead [52] ResNet50 27.866G 38.890M 0.249 0.335 0.189 0.328
-DDOD [53] ResNet50 46.514G 32.378M 0.253 0.335 0.230 0.351
-RTMDet [54] 51.278G 52.316M 0.229 0.349 0.212 0.350
-EfficientDet [55] 34.686G 18.320M 0.146 0.464 0.094 0.517
+${I}_{\text{topk}}\in\mathbb{N}^{K}$ represents the corresponding channel indices for these weights.
+This dynamic selection mechanism enables adaptive identification of the most informative channels for each sample, thereby enhancing the effectiveness of cross-task feature interaction.
+TABLE II: Comparison with Other State-of-the-art object detection methods on DenseSIRST.
+Method Backbone FLOPs Params DenseSIRST
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+                         End2End
+DETR [24] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000
+Sparse R-CNN [25] ResNet50 45.274G 0.106G 0.183 0.572 0.154 0.614
+Deformable DETR [26] ResNet50 51.772G 40.099M 0.024 0.016 0.018 0.197
+Conditional DETR [27] ResNet50 27.143G 40.297M 0.000 0.000 0.000 0.001
+DAB-DETR [28] ResNet50 28.939G 43.702M 0.005 0.054 0.000 0.001
 Two-stage
@@
 Dynamic R-CNN [35] ResNet50 63.179G 41.348M 0.184 0.235 0.111 0.190
+One-stage
+SSD [36] VGG16 87.552G 23.746M 0.211 0.421 0.178 0.424
+RetinaNet [37] ResNet50 52.203G 36.330M 0.114 0.510 0.086 0.523
+YOLOv3 [38] Darknet 50.002G 61.949M 0.233 0.424 0.207 0.413
+CenterNet [39] ResNet50 50.278G 32.111M 0.138 0.316 0.124 0.317
+FCOS [22] ResNet50 50.291G 32.113M 0.232 0.315 0.204 0.324
+ATSS [40] ResNet50 51.504G 32.113M 0.248 0.327 0.202 0.326
+CentripetalNet [41] HourglassNet 0.491T 0.206G 0.244 0.259 0.201 0.244
+AutoAssign [42] ResNet50 50.555G 36.244M 0.255 0.354 0.180 0.314
+GFL [43] ResNet50 52.296G 32.258M 0.264 0.367 0.230 0.317
+PAA [44] ResNet50 51.504G 32.113M 0.255 0.545 0.228 0.551
+VFNet [45] ResNet50 48.317G 32.709M 0.253 0.336 0.214 0.336
+PVT-T [46] PVT 41.623G 21.325M 0.109 0.481 0.093 0.501
+YOLOF [47] ResNet50 25.076G 42.339M 0.091 0.009 0.002 0.009
+YOLOX [48] CSPDarknet 8.578G 8.968M 0.210 0.341 0.180 0.331
+TOOD [7] ResNet50 50.456G 32.018M 0.256 0.355 0.226 0.342
+DyHead [49] ResNet50 27.866G 38.890M 0.249 0.335 0.189 0.328
+DDOD [50] ResNet50 46.514G 32.378M 0.253 0.335 0.230 0.351
+RTMDet [51] CSPNeXt 51.278G 52.316M 0.229 0.349 0.212 0.350
+EfficientDet [52] EfficientNet 34.686G 18.320M 0.146 0.464 0.094 0.517
+$\star$ BAFE-Net (Ours) ResNet50 71.639G 35.626M 0.274 0.342 0.248 0.338
+IV-B2 Cross-Task Feature Hard-Exchange
+The introduction of feature hard-exchange is fundamentally driven by the limitations of conventional feature fusion methods in infrared small target detection.
+Soft fusion techniques such as element-wise addition or concatenation often compromise discriminative power by diluting sparse target features through weighted averaging, while simultaneously permitting gradient conflicts between detection and segmentation objectives during backpropagation.
+Furthermore, the inherent semantic gap between heterogeneous tasks—object detection versus background segmentation—causes modality mismatch when transferring unadapted features directly.
+Feature hard-exchange directly addresses these challenges by enabling replacement of the most discriminative channels identified by DCS Module, ensuring explicit semantic transfer without feature dilution, maintaining gradient isolation through binary masking, and resolving cross-modal discrepancies via feature adaptation.
+The cross-task feature hard-exchange mechanism operates through three sequential steps.
+  • Exchange Mask Generation: To perform feature exchange, let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ be the input tensors from the target detection and background segmentation tasks, respectively, each with a shape of $(C,H,W)$, where $C$ is the number of channels, and $H$ and $W$ are the height and width.
+Using channel indices $I_{\text{topk}}$ from the DCS module (Eq.
+8), we construct a binary exchange mask $\mathbf{e}\in\mathbb{R}^{C}$:
+    $\mathbf{e}[i]=\begin{cases}1,&\mathrm{if}\ i\in I_{\text{topk}}\\ 0,&\mathrm{otherwise}\end{cases},$ (9)
+    where $i$ denotes channel index. This mask $\mathbf{e}$ identifies feature channels for exchange. The 1D channel mask is broadcast to spatial dimensions:
+    $\mathbf{E}_{d}=\mathbf{e}\otimes\mathbf{1}_{H\times W},$ (10)
+    where $\otimes$ indicates broadcasting operation, and $\mathbf{1}_{H\times W}$ is an all-one matrix.
+Detection and segmentation branches generate $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ respectively.
+  • Feature Exchange and Adaptation: Bidirectional hard exchange is performed as:
+    $\mathbf{Y}_{1}=\mathbf{X}_{1}\odot(\mathbf{1}-\mathbf{E}_{d1})+\mathcal{A}\left(\mathbf{X}_{2}\odot\mathbf{E}_{d2}\right),$ (11)
+    $\mathbf{Y}_{2}=\mathbf{X}_{2}\odot(\mathbf{1}-\mathbf{E}_{d2})+\mathcal{A}\left(\mathbf{X}_{1}\odot\mathbf{E}_{d1}\right),$ (12)
+    where $\odot$ denotes element-wise multiplication, $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ represent the output features, and $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ are the dynamic masks.
+To address the cross-modal nature of features, we introduce an adapter module $\mathcal{A}$, implemented as a multi-layer perceptron (MLP), which projects features into a common space before exchange.
+This design enables effective cross-modal interaction while preserving modality-specific characteristics.
+Compared to soft fusion techniques, hard-exchange provides distinct advantages: complete channel replacement preserves feature integrity by avoiding partial feature loss; gradient isolation ensures that gradients propagate exclusively to the source task for exchanged channels; and binary masks enable computationally efficient zero-parameter exchange.
+This integrated design maintains high discriminative power in complex infrared scenarios while ensuring feature compatibility through the adapter $\mathcal{A}$.
+TABLE III: Comparison with Other State-of-the-art object detection methods on IRSTD-1k and IRSDSS.
+Method Backbone FLOPs Params IRSTD-1k IRSDSS
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
                         End2End
-DETR [63] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000
-Sparse R-CNN [64] ResNet50 45.274G 0.106G 0.183 0.572 0.154 0.614
-Deformable DETR [65] ResNet50 51.772G 40.099M 0.024 0.016 0.018 0.197
-Conditional DETR [66] ResNet50 27.143G 40.297M 0.000 0.000 0.000 0.001
-DAB-DETR [67] ResNet50 28.939G 43.702M 0.005 0.054 0.000 0.001
-$\star$ BAFE-Net (Ours) ResNet18 57.654G 22.31M 0.283 0.335 0.233 0.325
-$\star$ BAFE-Net (Ours) ResNet50 71.639G 35.626M 0.274 0.342 0.248 0.338
-Figure 6: Feature visualization comparison. Compared with other methods, the proposed BAFE-Net shows superior performance with higher detection of small objects and fewer false alarms.
-V-B Comparison with State-of-the-Arts
-We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall.
-The experimental results are presented in Table II.
-End-to-end (End2End) methods, while theoretically appealing due to their NMS-free style, demonstrated suboptimal performance.
-This can be attributed to their inherent requirement for extensive training data, akin to COCO-level volumes, for effective operation.
-The relatively limited availability of infrared small target detection datasets inherently handicaps such methods, resulting in their diminished performance.
-Next, we observed that two-stage methods underperformed compared to one-stage methods.
-Two-stage methods generate candidate regions based on deeper features, which are then subjected to classification and bounding box regression.
-However, the DenseSIRST dataset, characterized by small and densely clustered infrared targets, poses a challenge for such methods, leading to their suboptimal performance.
-In contrast, one-stage methods, which directly regress bounding boxes and categories on the feature pyramids, demonstrated superior performance.
-These methods are not only faster and more computationally efficient, but also achieved higher mAP and recall, indicating their overall effectiveness.
-Significantly, our BAFE-Net method outperformed all other one-stage methods.
-The superior performance of BAFE-Net validates the efficacy of our approach in accurately detecting small infrared targets amidst dense conditions.
-TABLE III: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net
-Strategy Module DenseSIRST
-Copy-Paste Cluster Gaussian Background-Aware mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) 0.225 0.311 0.180 0.314
-(b) $\checkmark$ 0.236 0.322 0.203 0.303
-(c) $\checkmark$ $\checkmark$ 0.242 0.329 0.205 0.335
-(d) $\checkmark$ $\checkmark$ 0.245 0.321 0.211 0.338
-(e) $\checkmark$ $\checkmark$ $\checkmark$ 0.250 0.319 0.215 0.315
-(f) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.258 0.340 0.220 0.340
-TABLE IV: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
+DETR [24] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000 0.000 0.001 0.000 0.001
+Sparse R-CNN [25] ResNet50 45.274G 0.106G 0.136 0.370 0.160 0.388 0.826 0.953 0.827 0.952
+Deformable DETR [26] ResNet50 51.772G 40.099M 0.091 0.018 0.015 0.126 0.551 0.782 0.550 0.770
+Conditional DETR [27] ResNet50 27.143G 40.297M 0.000 0.001 0.000 0.002 0.000 0.010 0.000 0.011
+DAB-DETR [28] ResNet50 28.939G 43.702M 0.002 0.002 0.000 0.005 0.346 0.769 0.219 0.762
+Two-stage
+Faster R-CNN [29] ResNet50 0.759T 33.035M 0.298 0.346 0.269 0.344 0.416 0.401 0.478 0.524
+Cascade R-CNN [30] ResNet50 90.978G 69.152M 0.325 0.370 0.323 0.377 0.780 0.818 0.765 0.803
+Grid R-CNN [31] ResNet50 0.177T 64.467M 0.299 0.402 0.306 0.393 0.746 0.826 0.720 0.786
+Libra R-CNN [32] ResNet50 63.990G 41.611M 0.323 0.395 0.291 0.351 0.771 0.805 0.711 0.772
+TridentNet [33] ResNet50 0.759T 33.035M 0.311 0.341 0.288 0.352 0.672 0.757 0.631 0.701
+SABL [34] ResNet50 0.125T 42.213M 0.298 0.340 0.296 0.401 0.681 0.750 0.730 0.812
+Dynamic R-CNN [35] ResNet50 63.179G 41.348M 0.314 0.355 0.309 0.372 0.761 0.805 0.659 0.738
+One-stage
+SSD [36] VGG16 87.552G 23.746M 0.296 0.427 0.279 0.480 0.847 0.919 0.842 0.947
+RetinaNet [37] ResNet50 52.203G 36.330M 0.368 0.564 0.298 0.554 0.825 0.933 0.842 0.936
+YOLOv3 [38] Darknet 50.002G 61.949M 0.200 0.311 0.158 0.342 0.779 0.918 0.830 0.881
+CenterNet [39] ResNet50 50.278G 32.111M 0.201 0.497 0.202 0.544 0.853 0.924 0.847 0.916
+FCOS [22] ResNet50 50.291G 32.113M 0.307 0.426 0.303 0.446 0.841 0.907 0.842 0.924
+ATSS [40] ResNet50 51.504G 32.113M 0.355 0.429 0.355 0.475 0.849 0.911 0.861 0.907
+CentripetalNet [41] HourglassNet 0.491T 0.206G 0.232 0.333 0.245 0.375 0.850 0.918 0.855 0.911
+AutoAssign [42] ResNet50 50.555G 36.244M 0.326 0.416 0.336 0.433 0.842 0.918 0.851 0.925
+GFL [43] ResNet50 52.296G 32.258M 0.269 0.337 0.356 0.478 0.857 0.917 0.864 0.924
+PAA [44] ResNet50 51.504G 32.113M 0.356 0.668 0.342 0.573 0.957 0.941 0.868 0.933
+VFNet [45] ResNet50 48.317G 32.709M 0.360 0.427 0.362 0.456 0.848 0.908 0.854 0.912
+PVT-T [46] PVT 41.623G 21.325M 0.360 0.546 0.360 0.503 0.843 0.931 0.855 0.933
+YOLOF [47] ResNet50 25.076G 42.339M 0.216 0.304 0.146 0.276 0.341 0.441 0.281 0.429
+YOLOX [48] CSPDarknet 8.578G 8.968M 0.367 0.533 0.316 0.476 0.850 0.917 0.844 0.911
+TOOD [7] ResNet50 50.456G 32.018M 0.297 0.390 0.286 0.427 0.860 0.914 0.864 0.908
+DyHead [49] ResNet50 27.866G 38.890M 0.350 0.431 0.296 0.389 0.846 0.925 0.849 0.923
+DDOD [50] ResNet50 46.514G 32.378M 0.281 0.362 0.338 0.456 0.855 0.921 0.856 0.928
+RTMDet [51] CSPNeXt 51.278G 52.316M 0.276 0.476 0.216 0.450 0.332 0.765 0.353 0.760
+EfficientDet [52] EfficientNet 34.686G 18.320M 0.279 0.442 0.359 0.560 0.849 0.934 0.829 0.911
+$\star$ BAFE-Net (Ours) ResNet50 71.639G 35.626M 0.410 0.545 0.370 0.497 0.860 0.900 0.868 0.938
+TABLE IV: Comparison with Other State-of-the-art infrared target detection methods on DenseSIRST, IRSTD-1k and IRSDSS.
+Method DenseSIRST IRSTD-1k IRSDSS
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  ACM [11] 0.116 0.163 0.115 0.060 0.187 0.307 0.146 0.210 0.475 0.560 0.460 0.452
+DNANet [13] 0.188 0.219 0.135 0.139 0.161 0.245 0.196 0.220 0.705 0.807 0.786 0.789
+AGPCNet [14] 0.154 0.183 0.171 0.139 0.222 0.268 0.212 0.255 0.532 0.645 0.560 0.590
+UIUNet [53] 0.178 0.235 0.208 0.198 0.349 0.475 0.192 0.175 0.771 0.761 0.831 0.826
+MTUNet [54] 0.272 0.337 0.197 0.152 0.125 0.289 0.242 0.344 0.627 0.737 0.709 0.707
+RDIAN [55] 0.101 0.124 0.172 0.128 0.107 0.149 0.115 0.138 0.650 0.754 0.694 0.731
+URANet [56] 0.201 0.224 0.185 0.155 0.181 0.570 0.200 0.571 0.694 0.794 0.807 0.832
+ABCNet [57] 0.261 0.315 0.189 0.159 0.335 0.406 0.202 0.196 0.684 0.842 0.763 0.786
+SeRankDet [58] 0.216 0.261 0.114 0.120 0.308 0.408 0.160 0.268 0.658 0.742 0.774 0.788
+$\star$ BAFE-Net (Ours) 0.274 0.342 0.248 0.338 0.410 0.545 0.370 0.497 0.860 0.900 0.868 0.938
+V Experiments
+V-A Experimental Settings
+1) Dataset: During the experimental phase of our research, we conducted tests on the proposed dataset DenseSIRST.
+In addition to our proposed dataset, we also employed the publicly available IRSTD-1k dataset to further validate the effectiveness and generalization capability of our method.
+2) Implementation Details: Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion.
+We choose Focal Loss as the loss function of the classification branch, IoU Loss as the loss function of the regression branch, and Cross Entropy Loss as the loss function of the segmentation branch, with equal weighting assigned to three losses to ensure balanced multi-task optimization.
+We employ DAdaptAdam, a parameter-free adaptive optimization algorithm that dynamically adjusts learning rates based on gradient statistics, thereby improving training stability and reducing the need for manual learning rate tuning.
+The base learning rate is set to 1.0, and the weight decay is set to 0.05.
+To balance training time and convergence, we train all models for 24 epochs within the DeepInfrared framework.
+3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment.
+TABLE V: Ablation study on the impact of Segmentation Head, Feature Exchange and Adapter on detection performance in Background-Aware Feature Exchange
 Strategy Module DenseSIRST
@@
 (d) $\checkmark$ $\checkmark$ $\checkmark$ 0.262 0.318 0.208 0.305
-TABLE V: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
+TABLE VI: Ablation study on the impact of Hidden Layers on detection performance in Adapter
 Strategy Hidden Layers DenseSIRST
 mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) 8 0.222 0.305 0.206 0.308
-(b) 16 0.243 0.318 0.214 0.310
-(c) 32 0.256 0.329 0.218 0.312
-(d) 64 0.266 0.336 0.230 0.323
-TABLE VI: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
+  (a) 8 0.243 0.327 0.206 0.308
+(b) 16 0.252 0.337 0.214 0.310
+(c) 32 0.254 0.326 0.218 0.312
+(d) 64 0.260 0.329 0.230 0.323
+(e) 128 0.254 0.326 0.211 0.333
+(f) 256 0.243 0.319 0.212 0.312
+TABLE VII: Ablation study on the impact of Exchange Method on detection performance in Feature Exchange Mechanism
 Strategy Exchange Method DenseSIRST
@@
 (d) Spatial + Channel 0.224 0.310 0.201 0.328
-TABLE VII: Ablation study on the impact of feature selection on detection performance in Channel-wise Feature Exchange
+TABLE VIII: Ablation study on the impact of feature selection on detection performance in Channel-wise Feature Exchange
 Strategy Feature Selection DenseSIRST
@@
 (c) Dynamic Selection 0.283 0.335 0.233 0.325
-TABLE VIII: Ablation study on the impact of the number of exchange channels on detection performance in Top-K selection
+TABLE IX: Ablation study on the impact of the number of exchange channels on detection performance in Top-K selection
 Strategy P DenseSIRST
@@
 (g) 7/8 0.238 0.316 0.214 0.329
-(h) 1/16 0.245 0.326 0.213 0.301
-(i) 1/32 0.236 0.319 0.212 0.312
-TABLE IX: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
-Strategy ResNet DenseSIRST
-mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
-  (a) ResNet18 0.283 0.335 0.233 0.341
-(b) ResNet34 0.281 0.354 0.226 0.332
-(c) ResNet50 0.274 0.342 0.248 0.338
-(d) ResNet101 0.261 0.317 0.241 0.336
+TABLE X: Ablation study on the impact of different ResNet on detection performance in BAFE-Net
+Strategy ResNet DenseSIRST IRSTD-1k
+mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
+  (a) ResNet18 0.283 0.335 0.233 0.325 0.433 0.603 0.380 0.528
+(b) ResNet34 0.281 0.354 0.226 0.332 0.385 0.495 0.381 0.564
+(c) ResNet50 0.274 0.342 0.248 0.338 0.410 0.545 0.370 0.497
+(d) ResNet101 0.261 0.317 0.241 0.336 0.385 0.520 0.378 0.561
+V-B Comparison with State-of-the-Arts
+We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall.
+The experimental results for general object detection frameworks are presented in Table II.
+End-to-end (End2End) methods, while theoretically appealing due to their NMS-free style, demonstrated suboptimal performance on both datasets.
+This can be attributed to their inherent requirement for extensive training data, akin to COCO-level volumes, for effective operation.
+The relatively limited availability of infrared small target detection datasets inherently handicaps such methods, resulting in their diminished performance.
+Next, we observed that two-stage methods underperformed compared to one-stage methods.
+Two-stage methods generate candidate regions based on deeper features, which are then subjected to classification and bounding box regression.
+However, the DenseSIRST dataset, characterized by small and densely clustered infrared targets, poses a challenge for such methods, leading to their suboptimal performance.
+In contrast, one-stage methods, which directly regress bounding boxes and categories on the feature pyramids, demonstrated superior performance on both datasets.
+These methods are not only faster and more computationally efficient, but also achieved higher mAP and recall, indicating their overall effectiveness.
+Significantly, our BAFE-Net method outperformed all other one-stage methods on both datasets.
+Beyond the DenseSIRST dataset, we further conducted experiments on the IRSTD-1k[5] and IRSDSS[59] datasets, as summarized in Table III.
+The results on IRSTD-1k were consistent with those observed on DenseSIRST, again highlighting the advantage of BAFE-Net in handling dense small-target detection.
+In contrast, the IRSDSS dataset mainly consists of maritime targets with relatively large object sizes.
+Consequently, while BAFE-Net still ranked among the best-performing approaches, the differences among the top methods on this dataset were relatively minor.
+Notably, BAFE-Net achieved state-of-the-art performance, with mAP scores matching or surpassing those of the best competing methods.
+The superior performance of BAFE-Net consistently validates the efficacy of our approach in accurately detecting small infrared targets under diverse detection scenarios, including both densely clustered and sparsely distributed configurations.
+Furthermore, to comprehensively situate BAFE-Net within the specific domain of infrared small target detection (IRSTD), we conducted comparisons against state-of-the-art methods specifically designed for this task.
+The results of these comparisons on both DenseSIRST, IRSTD-1k and IRSDSS datasets are presented in Table IV.
+As shown, BAFE-Net also achieves superior performance compared to these specialized IRSTD methods, further demonstrating its effectiveness.
 V-C Visual Analysis
+Figure 5: Feature visualization comparison.
+This comparison showcases the detection results of different methods in terms of feature visualization.
+The proposed BAFE-Net demonstrates superior performance by effectively detecting small objects with fewer false alarms.
+V-D Ablation Study
+V-D1 Ablation Study on the Scheme of Background-Aware Feature Exchange
+The BAFE module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table V.
+To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
+Table V summarizes the results and reveals several key insights.
+First, each individual module contributes significantly and independently to detection performance, thereby validating the foundational design of our framework.
+Second, a comparison between Strategy (a) and Strategy (b) highlights the benefit of incorporating background-based segmentation into the object detection pipeline, where the detection head interacts with the segmentation head.
+This integration results in performance gains of 2.1% and 0.3% on the DenseSIRST dataset.
+Furthermore, the performance difference between Strategy (b) and Strategy (c) demonstrates the effectiveness of the Feature Exchange mechanism in enhancing the interaction between the detection and segmentation heads.
+This enhancement leads to improvements of 0.8% and 1.6% on the DenseSIRST dataset.
+Finally, comparing Strategy (c) with Strategy (d) shows that introducing a feature adapter prior to the Feature Exchange further amplifies detection performance.
+As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration.
+Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity.
+To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table VI.
+The results show that performance consistently improves as the number of hidden layers increases, up to 64 layers.
+However, when the number of hidden layers exceeds 64, the detection performance begins to decline, while the parameter count continues to grow significantly.
+Therefore, we select 64 hidden layers for the Adapter, as it offers the best trade-off between detection accuracy and computational overhead.
+V-D2 Ablation Study on Feature Exchange Mechanism
+In this section, we conduct a rigorous investigation of various Feature Exchange Mechanisms, including Channel Exchange, Spatial Exchange, and several of their combinations.
+The goal is to identify the most effective strategy for facilitating cross-task feature interaction.
+Channel Exchange enables direct information flow between features at the channel level, while Spatial Exchange operates by exchanging information along the spatial dimensions of the feature maps.
+To ensure an equitable exchange of features, the exchange ratio is consistently fixed at 0.5 in all configurations.
+As shown in Table VII, Channel Exchange consistently outperforms other Feature Exchange strategies.
+The experimental results clearly highlight its effectiveness and stability, confirming its selection as the optimal mechanism for feature exchange within our model.
+V-D3 Ablation Study in Dynamic Channel Selection
+In this section, we perform an in-depth investigation of different feature selection strategies, including random selection, fixed selection, and dynamic selection, with the goal of identifying the most effective approach for enhancing feature exchange.
+1) Randomly Selection: This method selects feature channels randomly, which may lead to suboptimal feature exchanges.
+In some cases, it could even select features that hinder the task, ultimately reducing model performance.
+2) Fixed Selection: While fixed selection provides consistency during training and eliminates uncertainty, it lacks the adaptability to dynamic input features, thus limiting the model’s performance in complex or varied scenarios.
+3) Dynamic Selection: dynamic selection maximizes the performance of the model and reduces unnecessary information redundancy by selecting the most useful feature channels for exchange.
+The exchange ratio is judiciously set to 0.5 to maintain a balance.
+As shown in Table VIII, the advantages of dynamic selection are obvious.
+Empirical evidence highlights the robustness and optimality of dynamic selection, thus confirming its adoption as the preferred method for feature selection in our model.
+Additionally, we conduct a focused analysis on the impact of the hyperparameter $p$ in the Top-K selection mechanism.
+As shown in Table IX, setting $p=1/2$ yields the best performance, validating this choice as optimal for our design.
+V-D4 Impact of Backbone Depth
+In infrared small target detection, the depth of the backbone network plays a pivotal role in determining detection performance.
+A shallower backbone, though computationally efficient, may fail to capture sufficiently rich semantic features, thereby limiting the model’s ability to distinguish small targets from complex backgrounds.
+On the other hand, an overly deep backbone may introduce excessive abstraction, potentially suppressing the subtle features of small targets and leading to performance degradation.
+Therefore, achieving an optimal balance between depth and feature representation is essential for effective infrared target detection.
+We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net.
+The results of these experiments are summarized in Table X.
+Notably, on DenseSIRST, ResNet18 has the highest $mAP_{07}$ metric and ResNet50 has the highest $mAP_{12}$ metric; on IRSTD-1k, ResNet18 has the highest $mAP_{07}$ metric and ResNet34 has the highest $mAP_{12}$ metric.
 To further validate the effectiveness of our proposed methods, we conduct a visual analysis by comparing the feature activation maps of our approach with those of the top-performing competing methods in the quantitative evaluation.
@@
 The deeper activation levels in the target areas suggest that our method is more effective in capturing the discriminative features of the targets, contributing to improved detection accuracy.
-Second, our method generates cleaner background activation maps compared to the other methods.
+Second, our method generates cleaner background activation maps.
 As evident from Fig.
-6, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference.
+5, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference.
 This can be attributed to the background-aware mechanism employed by BAFE-Net, which learns to differentiate between target and background regions during the feature exchange process.
 By minimizing the activation levels in the background areas, our method demonstrates enhanced robustness against false positives, leading to a higher detection precision.
-V-D Ablation Study
-V-D1 Ablation Study on Background-Aware Gaussian Copy-Paste
-We initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting.
-The experimental outcomes, as outlined in Table III, generate several key insights, providing a holistic understanding of the impact of each module.
-Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study.
-The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
-Further refinement through clustering of pasted objects also yields substantial improvements.
-This is evident when comparing strategies (b) to (c), and (d) to (e), respectively, indicating the considerable value of object clustering.
-Applying Gaussian blur to the edges of the pasted objects, as a means to enhance their realistic appearance, also proves to be significantly beneficial in improving detection performance.
-This is substantiated when comparing strategies (b) with (d), and (c) with (e).
-Finally, the culmination of all modules in a Background-Aware Gaussian Copy-Paste context, as depicted by strategy (f), results in further enhancements in detection capability.
-This is demonstrated with an additional increase of 0.8% on the DenseSIRST dataset when comparing strategy (e) with strategy (f).
-V-D2 Ablation Study on the Scheme of Background-Aware Feature Exchange
-The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table IV.
-To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
-Table IV encapsulates the results of our experimental study and yields several crucial insights.
-Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research.
-Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head.
-This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
-Further, juxtaposition of Strategy (b) and Strategy (c) highlights the contribution of the Feature Exchange in the interaction mode between the detection head and the segmentation head.
-This interaction led to a notable improvement in detection performance by 1.1% on the DenseSIRST dataset.
-Upon comparing Strategy (c) and Strategy (d), it becomes evident that introducing feature adapation prior to the Feature Exchange further amplifies detection performance.
-As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration.
-Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity.
-To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  V.
-The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance.
-However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially.
-Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
-V-D3 Ablation Study on Feature Exchange Mechanism
-In this part, a rigorous exploration of various Feature Exchange Mechanisms was undertaken.
-This encompassed diverse mechanisms such as Channel Exchange and Spatial Exchange, along with a multitude of their combinations.
-The quest was to ascertain the most efficacious method for Feature Exchange.
-Channel Exchange, as a mechanism, facilitates the direct interplay at the channel level amongst different features.
-In contrast, Spatial Exchange is implemented by orchestrating exchanges along the spatial dimensions inherent in the feature maps.
-The exchange ratio was judiciously set at 0.5 to maintain a balance.
-As delineated in Table VI, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested.
-The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
-V-D4 Ablation Study in Dynamic Channel Selection
-In this section, we first conduct a rigorous exploration of feature selection methods.
-This includes random selection, fixed selection, and dynamic selection.
-The task is to determine the most effective feature selection method.
-1) Randomly Selection: This method selects feature channels randomly, which may lead to suboptimal feature exchanges.
-In some cases, it could even select features that hinder the task, ultimately reducing model performance.
-2) Fixed Selection: While fixed selection provides consistency during training and eliminates uncertainty, it lacks the adaptability to dynamic input features, thus limiting the model’s performance in complex or varied scenarios.
-3) Dynamic Selection: dynamic selection maximizes the performance of the model and reduces unnecessary information redundancy by selecting the most useful feature channels for exchange.
-The exchange ratio is judiciously set to 0.5 to maintain a balance.
-As shown in Table VII, the advantages of dynamic selection are obvious.
-Empirical evidence highlights the robustness and optimality of dynamic selection, thus confirming its adoption as the preferred method for feature selection in our model.
-In addition, we conduct a rigorous exploration of the value of $p$ in Top-K selection. As shown in Table VIII, $p$ is optimal when it is 1/2.
-V-D5 Impact of Backbone Depth
-For infrared small target detection, the depth of the backbone network emerges as a critical factor.
-A shallower network, while computationally efficient, might extract features with insufficient semantic richness, thereby impinging on the overall detection performance.
-Conversely, an excessively deep network might overwhelm the delicate features of smaller targets amidst the background, consequently leading to a performance degradation.
-We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net.
-The results of these experiments are summarized in Table IX.
-Notably, ResNet18 has the highest $mAP_{07}$ metric and ResNet50 has the highest $mAP_{12}$ metric.
 VI Conclusion
@@
 We introduce DenseSIRST, a benchmark dataset with detailed background annotations, enabling the transition from sparse to dense target detection.
-Building upon this, we propose BAFE-Net, a multi-task architecture that explicitly models background semantics and facilitates cross-task feature hard-exchange.
-Moreover, we develop BAG-CP, a data augmentation method that selectively pastes targets into sky regions, outperforming the original Copy-Paste method.
-Extensive experiments validate the effectiveness of BAFE-Net and BAG-CP in improving detection accuracy and reducing false alarms.
+Building upon this, we propose BAFE-Net, a multi-task architecture that explicitly models background semantics and facilitates cross-task feature hard-exchange, promoting the exchange of information between the classification and background segmentation branches.
+Moreover, we develop BAG-CP, a synthesis method for the DenseSIRST dataset, which also serves as a data augmentation technique.
+BAG-CP selectively pastes targets into sky regions, outperforming the original Copy-Paste method.
+Extensive experiments validate the effectiveness of BAFE-Net in improving detection accuracy and reducing false alarms.
 Our work underscores the importance of explicitly leveraging contextual information in infrared small target detection and paves the way for the development of more robust and context-aware algorithms in this domain.